### v1

# Likelihood-Ratio Regularized Quantile Regression: Adapting Conformal Prediction to High-Dimensional Covariate Shifts

Abstract
We consider the problem of conformal prediction under covariate shift. Given labeled data from a source domain and unlabeled data from a covariate shifted target domain, we seek to construct prediction sets with valid marginal coverage in the target domain. Most existing methods require estimating the unknown likelihood ratio function, which can be prohibitive for high-dimensional data such as images. To address this challenge, we introduce the likelihood ratio regularized quantile regression (LR-QR) algorithm, which combines the pinball loss with a novel choice of regularization in order to construct a threshold function without directly estimating the unknown likelihood ratio. We show that the LR-QR method has coverage at the desired level in the target domain, up to a small error term that we can control. Our proofs draw on a novel analysis of coverage via stability bounds from learning theory. Our experiments demonstrate that the LR-QR algorithm outperforms existing methods on high-dimensional prediction tasks, including a regression task for the Communities and Crime dataset, and an image classification task from the WILDS repository.
1 Introduction
Conformal prediction is a framework to construct distribution-free prediction sets for black-box predictive models (e.g., Saunders et al., , 1999; Vovk et al., , 1999, 2022, etc). Formally, given a pretrained prediction model $f:\mathcal{X}\to\mathcal{Y}$ that maps features $x\in\mathcal{X}$ to labels $y\in\mathcal{Y}$, as well as $n_{1}$ calibration datapoints $(X_{i},Y_{i}):i\in[n_{1}]$ sampled i.i.d. from a calibration distribution $\mathbb{P}_{1}$, we seek to construct a prediction set $C(X_{\textnormal{test}})\subseteq\mathcal{Y}$ for test features $X_{\textnormal{test}}$ sampled from a test distribution $\mathbb{P}_{2}$. We aim to cover the true label $Y_{\textnormal{test}}$ with probability at least $1-\alpha$ for some $\alpha\in(0,1)$: that is, $\mathbb{P}(Y_{\textnormal{test}}\in C(X_{\textnormal{test}}))\geqslant 1-\alpha$. The left-hand side of this inequality is the marginal coverage of the prediction set $C$, averaged over the randomness of both the calibration datapoints and the test datapoint. In the case that the calibration and test distributions coincide ($\mathbb{P}_{1}=\mathbb{P}_{2}$), there are numerous conformal prediction algorithms that construct distribution-free prediction sets with valid marginal coverage; for instance, split and full conformal prediction (e.g., Papadopoulos et al., , 2002; Lei et al., , 2013).
However, in practice, it is often the case that test data is sampled from a different distribution than calibration data. This general phenomenon is known as distribution shift (e.g., Quiñonero-Candela et al., , 2009; Sugiyama and Kawanabe, , 2012). One particularly common type of distribution shift is covariate shift (Shimodaira, , 2000), where the conditional distribution of $Y|X$ stays fixed, but the marginal distribution of features changes from calibration to test time. For instance, in the setting of image classification for autonomous vehicles, the calibration and test data might have been collected under different weather conditions Yu et al., (2020); Koh et al., (2021). Under covariate shift, ordinary conformal prediction algorithms may lose coverage.
Recently, a number of methods have been proposed to adapt conformal prediction to covariate shift, e.g., in Tibshirani et al., (2019); Park et al., 2022b ; Park et al., 2022a ; Gibbs et al., (2023); Qiu et al., (2023); Yang et al., (2024); Gui et al., (2024). Most existing approaches attempt to estimate the likelihood ratio function $r:\mathcal{X}\to\mathbb{R}$, defined as $r(x)=(d\mathbb{P}_{2,X}/d\mathbb{P}_{1,X})(x)$, for all $x\in\mathcal{X}$. One can construct an estimate $\hat{r}$ of the likelihood ratio if one has access to additional unlabeled datapoints sampled i.i.d from the test distribution $\mathbb{P}_{2}$. Methods for likelihood ratio estimation include using Bayes’ rule to express it as a ratio of classifiers Friedman, (2003); Qiu et al., (2023) and domain adaptation Ganin and Lempitsky, (2015); Park et al., 2022b . However, such estimates may be inaccurate for high-dimensional data. This error propagates to the coverage of the resulting conformal predictor, and the prediction sets may no longer attain the nominal coverage level. Thus, it is natural to ask the following question:
Can one design a conformal prediction algorithm that attains valid coverage in the target domain, without estimating the entire function $r$?
In this paper, we present a method that answers this question in the affirmative. We construct our prediction sets by introducing and solving a regularized quantile regression problem, which combines the pinball loss with a novel data-dependent regularization term that can be computed from one-dimensional projections of the likelihood ratio $r$. Crucially, the objective function can be estimated at the parametric rate, with only a mild dependence on the dimension of the feature space. This regularization is specifically chosen to ensure that the first order conditions of the pinball loss lead to coverage at test-time. Geometrically, it turns out that the regularization aligns the selected threshold function with the true likelihood ratio $r$. The resulting method, which we call likelihood ratio regularized quantile regression (LR-QR), outperforms existing methods on high-dimensional datasets with covariate shift.
Our contributions include the following:
  • We propose the LR-QR algorithm, which constructs a conformal predictor that adapts to covariate shift without directly estimating the likelihood ratio.
  • We show that the minimizers of the population LR-QR objective have coverage in the test distribution. We also show that the minimizers of the empirical LR-QR objective lead to coverage up to a small error term that we can control, by drawing on a novel analysis of coverage via stability bounds from learning theory.
  • We demonstrate the effectiveness of the LR-QR algorithm on high-dimensional datasets under covariate shift, including the Communities and Crime dataset and the RxRx1 dataset from the WILDS repository. Here, we leverage our theoretical insights, by choosing the regularization parameter by a cross validation procedure guided by theory. An implementation of LR-QR can be accessed at the following link: https://github.com/shayankiyani98/LR-QR.
The structure of this paper is as follows. In Section 2, we rigorously state the problem. In Section 3, we present our method, as well as intuitions behind it. In Section 3.5, we present the algorithm. In Section 4 we present our theoretical results, in both the infinite sample and finite sample settings. In Section 5, we present our experimental results on high-dimensional datasets with covariate shift. All proofs are deferred to the appendix.
1.1 Related work
Here we only list prior work most closely related to our method; we provide more references in Appendix C. The early ideas of conformal prediction were developed in Saunders et al., (1999); Vovk et al., (1999). With the rise of machine learning, conformal prediction has emerged as a widely used framework for constructing prediction sets (e.g., Papadopoulos et al., , 2002; Vovk et al., , 2005; Vovk, , 2013). Classical conformal prediction guarantees validity when the calibration and test data are drawn from the same distribution. In contrast, when there is distribution shift between the calibration and test data (e.g., Quiñonero-Candela et al., , 2009; Shimodaira, , 2000; Sugiyama and Kawanabe, , 2012; Ben-David et al., , 2010; Taori et al., , 2020), coverage may not hold. Covariate shift is a type of dataset shift that arises in many settings, e.g., when predicting disease risk for individuals whose features may evolve over time, while the outcome distribution conditioned on the features remains stable (Quiñonero-Candela et al., , 2009).
Numerous works have addressed conformal prediction under various types of distribution shift (Tibshirani et al., , 2019; Park et al., 2022b, ; Park et al., 2022a, ; Qiu et al., , 2023; Si et al., , 2024). For example, Tibshirani et al., (2019) investigated conformal prediction under covariate shift, assuming the likelihood ratio between source and target covariates is known. Lei and Candès, (2021) allowed the likelihood ratio to be estimated, rather than assuming it is known. Park et al., 2022b developed prediction sets with a calibration-set conditional (PAC) property under covariate shift. Qiu et al., (2023); Yang et al., (2024) developed prediction sets with asymptotic coverage that are doubly robust in the sense that their coverage error is bounded by the product of the estimation errors of the quantile function of the score and the likelihood ratio. Cauchois et al., (2024) construct prediction sets based on a distributionally robust optimization approach.
In contrast, our algorithm entirely avoids estimating the likelihood ratio function. Rather, it works by constructing a novel regularized regression objective, whose stationary conditions ensure coverage in the test domain. We can minimize the objective by estimating certain expectations of the data distribution—which implicitly involve estimating only certain functionals of the likelihood ratio. We further show that the coverage is retained in finite samples via a novel analysis of coverage leveraging stability bounds Shalev-Shwartz et al., (2010); Shalev-Shwartz and Ben-David, (2014). We illustrate that our algorithms behave better in high-dimensional datasets than existing methods.
Aiming to achieve coverage under a predefined set of covariate shifts, Gibbs et al., (2023) develop an approach based on minimizing the quantile loss over a linear function class. We build on their approach, but develop a novel regularization scheme that allows us to effectively optimize over a data-driven class, adaptive to the unknown shift $r$.
2 Problem formulation
In this section we fix notation and state our problem.
2.1 Preliminaries and notations
For $\alpha\in(0,1)$, recall that the quantile (pinball) loss $\ell_{\alpha}$ is defined for all $c,s\in\mathbb{R}$ as
$\ell_{\alpha}(c,s):=\left\{\begin{array}[]{l}(1-\alpha)(s-c)\text{ if }s\geq c,\\ \alpha(c-s)\text{ if }s<c.\end{array}\right.$
For any distribution $P$, the minimizers of $c\mapsto\mathbb{E}_{S\sim P}[\ell_{\alpha}(c,S)]$ are the $(1-\alpha)$th quantiles of $P$.
Let the source or calibration distribution be denoted $\mathbb{P}_{1}=\mathbb{P}_{1,X}\times\mathbb{P}_{Y|X}$, and let the target or test distribution be denoted $\mathbb{P}_{2}=\mathbb{P}_{2,X}\times\mathbb{P}_{Y|X}$, a covariate shifted version of the calibration distribution. Let $\mathbb{E}_{i}$ denote the expectation over $\mathbb{P}_{i}$, $i=1,2$. Let $x\mapsto r(x)=\frac{d\mathbb{P}_{2,X}}{d\mathbb{P}_{1,X}}(x)$ denote the unknown likelihood ratio function.
Recall that a prediction set $C:\mathcal{X}\to 2^{\mathcal{Y}}$ has marginal $(1-\alpha)$-coverage in the test domain if $\mathbb{P}_{2}\left[Y\in C(X)\right]\geqslant 1-\alpha$. Observe that $\mathbb{P}_{2}\left[Y\in C(X)\right]$ can be rewritten as $\mathbb{E}_{2}[\mathbf{1}[Y\in C(X)]]$, where $\mathbf{1}[\cdot]$ denotes an indicator function. Let $S:(x,y)\mapsto S(x,y)$ denote the nonconformity score associated to a pair $(x,y)\in\mathcal{X}\times\mathcal{Y}$. Given a threshold function $q:\mathcal{X}\to\mathbb{R}$, we consider the corresponding conformal predictor $C:\mathcal{X}\to 2^{\mathcal{Y}}$ given by
$C(x)=\{y\in\mathcal{Y}:S(x,y)\leqslant q(x)\}$ (1)
for all $x\in\mathcal{X}$. Thus a threshold function $q$ yields a conformal predictor with marginal $(1-\alpha)$-coverage in the test domain if $\mathbb{P}_{2}[S(X,Y)\leqslant q(X)]\geqslant 1-\alpha$. We assume that $\alpha\leqslant 0.5$. For our theory, we consider $[0,1]$-valued scores.
[…]
2.2 Problem statement
We observe $n_{1}$ labeled calibration (or, source) datapoints $\{(X_{i},Y_{i}):i\in[n_{1}]\}$ drawn i.i.d. from the source distribution $\mathbb{P}_{1}$, and an additional $n_{3}$ unlabeled calibration datapoints $\mathcal{S}_{3}$. We also have $n_{2}$ unlabeled (target) datapoints $\mathcal{S}_{2}$ drawn i.i.d. from the target distribution $\mathbb{P}_{2}$. Given $\alpha\in(0,1)$, our goal is to construct a threshold function $q:\mathcal{X}\to\mathbb{R}$ that achieves marginal $(1-\alpha)$-coverage in the test domain: $\mathbb{P}_{2}[S(X,Y)\leqslant q(X)]\geqslant 1-\alpha.$
3 Algorithmic principles
Here we present some intuition behind our approach.
3.1 Change of measure
A key fact for us is the change-of-measure identity $\mathbb{E}_{1}[r(X)h(X)]=\mathbb{E}_{2}[h(X)]$ for all measurable functions $h:\mathcal{X}\to\mathbb{R}$. To see this, we simply write $\mathbb{E}_{1}[r(X)h(X)]=\mathbb{E}_{1}[d\mathbb{P}_{2,X}/d\mathbb{P}_{1,X}(X)\cdot h(X)]=\mathbb{E}_{2}[h(X)]$. Therefore, for a threshold function $q$,
$\mathbb{E}_{1}[r(X)\mathbf{1}[S(X,Y)\leqslant q(X)]]$ $=\mathbb{E}_{1}[r(X)\mathbb{P}_{Y|X}[S(X,Y)\leqslant q(X)]]$
$=\mathbb{E}_{2}[\mathbb{P}_{Y|X}[S(X,Y)\leqslant q(X)]]=\mathbb{P}_{2}[S(X,Y)\leqslant q(X)].$
Thus a threshold function $q$ yields a conformal predictor with marginal $(1-\alpha)$-coverage in the test domain if and only if
$\mathbb{E}_{1}[r(X)\mathbf{1}[S(X,Y)\leqslant q(X)]]\geqslant 1-\alpha.$
3.2 Warm-up: stationary conditions of pinball loss
As a warm-up, we begin with a simpler objective function, the pinball loss. Suppose that we knew the true likelihood ratio $r$. Then in order to obtain our threshold function, we could solve the optimization problem $\kappa^{*}=\arg\min_{\kappa\in\mathbb{R}}\mathbb{E}_{1}[\ell_{\alpha}(\kappa r(X),S)],$ and set our threshold function to equal $\zeta^{*}=\kappa^{*}r$. We now argue that the marginal $(1-\alpha)$-coverage of $\zeta^{*}$ in the test domain can be deduced from the first order condition with respect to $\kappa$. Assuming that the distribution of $S|X=x$ has a density for all $x\in\mathcal{X}$, and using the fact that $\frac{\partial}{\partial c}\ell_{\alpha}(c,s)=\mathbf{1}[s\leqslant c]-(1-\alpha)$ for all $c\neq s$, the derivative of the objective with respect to $\kappa$ is
$\frac{\partial}{\partial\kappa}\mathbb{E}_{1}[\ell_{\alpha}(\kappa r(X),S)]$ $=\mathbb{E}_{1}\left[\frac{\partial}{\partial\kappa}\ell_{\alpha}(\kappa r(X),S)\right]$
$=\mathbb{E}_{1}[r(X)(\mathbf{1}[S\leqslant\kappa r(X)]-(1-\alpha))].$
Setting this equal to zero, we deduce that the minimizer $\kappa^{*}$ satisfies $\mathbb{E}_{1}[r(X)(\mathbf{1}[S\leqslant\kappa^{*}r(X)]-(1-\alpha))]=0$. By the change-of-measure identity above, this implies the desired marginal coverage guarantee for $\zeta^{*}$.
The above problem can be recast as optimizing over the one-dimensional function class ~=$\tilde{\mathcal{H}}=\ \{\kappa r:\kappa\in\mathbb{R}\}$ spanned by $r$. So, if we knew $r$, we could ensure coverage by optimizing over a single scalar parameter, which would be convenient and also have good finite sample coverage. However, we do not know $r$. Therefore, we consider a general class $\mathcal{H}$ large enough to approximate $r$ in a suitable way, and solve an analogous optimization problem. Intriguingly, the first order conditions will ensure coverage.
Specifically, given a convex function class $\mathcal{H}$, we consider the optimization problem
$(\tilde{h}^{*},\kappa^{*})=\arg\min_{\tilde{h}\in\mathcal{H},\kappa\in\mathbb{R}}\mathbb{E}_{1}[\ell_{\alpha}(\kappa\tilde{h},S)],$
and set our threshold function equal to $\kappa^{*}\tilde{h}^{*}$. To derive the first order condition with respect to $\tilde{h}$, assume for simplicity that the minimizer $\tilde{h}^{*}$ lies in the interior of $\mathcal{H}$. Given a valid perturbation direction [^1] Formally, an element of the tangent cone of $\mathcal{H}$ at $\tilde{h}^{*}$. $g$, the function $\tilde{h}^{*}+\varepsilon g$ is in $\mathcal{H}$ for sufficiently small $\varepsilon$. Since for $\kappa=\kappa^{*}$, $\tilde{h}^{*}$ minimizes the objective in $\tilde{h}$, the function $\varepsilon\mapsto\mathbb{E}_{1}[\ell_{\alpha}(\kappa^{*}(\tilde{h}^{*}(X)+\varepsilon g(X)),S)]$ has a minimum at $\varepsilon=0$. Setting the derivative of this function with respect to $\varepsilon$ equal to zero as in Gibbs et al., (2023), we obtain the first order condition
$\frac{\partial}{\partial\varepsilon}\bigg{|}_{\varepsilon=0}\mathbb{E}_{1}[\ell_{\alpha}(\kappa^{*}(\tilde{h}^{*}(X)+\varepsilon g(X)),S)]$ $=\mathbb{E}_{1}[\kappa^{*}g(X)(\mathbf{1}[S\leqslant\kappa^{*}\tilde{h}^{*}(X)]-(1-\alpha))]$
$=0.$
If $\kappa^{*}\neq 0$ and $\mathbb{E}_{1}[g(X)]\neq 0$, then by rearranging and dividing by $\kappa^{*}\mathbb{E}_{1}[g(X)]$, this is equivalent to $\mathbb{E}_{1}\left[\frac{g(X)}{\mathbb{E}_{1}[g(X)]}\mathbf{1}[S\leqslant\kappa^{*}\tilde{h}^{*}(X)]\right]=1-\alpha$. If the likelihood ratio $r$ is a valid perturbation at $\tilde{h}^{*}$, we can plug in $g=r$ to find that the threshold function $\kappa^{*}\tilde{h}^{*}$ attains valid coverage in the test domain. Thus, we search for classes $\mathcal{H}$ such that $r$ is a valid perturbation at all interior points of $\mathcal{H}$.
3.3 Constrained quantile regression
We now design a suitable function class $\mathcal{H}$. Given $\varepsilon>0$, consider in this section $\mathcal{H}=B_{\varepsilon}(r)$, the ball of radius $\sqrt{\varepsilon}$ centered at $r$, consisting of all square-integrable functions $\tilde{h}$ satisfying $\mathbb{E}_{1}[(\tilde{h}(X)-r(X))^{2}]\leqslant\varepsilon$. Since the interior is an open set, all perturbations are valid. Thus, by our remarks above, if we minimize $\mathbb{E}_{1}[\ell_{\alpha}(\kappa\tilde{h}(X),S)]$ for scalar $\kappa$ and $\tilde{h}$ restricted to this ball, we obtain a threshold function with valid coverage under the covariate shift defined by $r$.
3.4 Removing explicit dependence on the likelihood ratio
Next, notice that we can expand $\mathbb{E}_{1}[(\tilde{h}(X)-r(X))^{2}]$ into
$\mathbb{E}_{1}[\tilde{h}(X)^{2}]+\mathbb{E}_{1}[-2r(X)\tilde{h}(X)]+\mathbb{E}_{1}[r(X)^{2}]=\mathbb{E}_{1}[\tilde{h}(X)^{2}]+\mathbb{E}_{2}[-2\tilde{h}(X)]+\mathbb{E}_{1}[r(X)^{2}],$
where we used the change-of-measure identity $\mathbb{E}_{1}[r(X)\tilde{h}(X)]=\mathbb{E}_{2}[\tilde{h}(X)]$. Only the first two terms depend on the optimization variable $\tilde{h}$, so it is enough to consider these two terms. This is where we make our second key observation. The insight is that neither of the terms $\mathbb{E}_{1}[\tilde{h}(X)^{2}]$ or $\mathbb{E}_{2}[-2\tilde{h}(X)]$ explicitly involve $r$, and thus they can be estimated by sample averages over the source and target data, respectively. Thus, we can minimize $\mathbb{E}_{1}[\ell_{\alpha}(\kappa\tilde{h}(X),S)]$ over $\kappa\in\mathbb{R}$ while keeping $\mathbb{E}_{1}[\tilde{h}(X)^{2}]+\mathbb{E}_{2}[-2\tilde{h}(X)]$ bounded. The threshold $\kappa^{*}\tilde{h}^{*}$ will have valid coverage under the covariate shift $r$.
3.5 Algorithm: likelihood ratio regularized quantile regression
Further, we can equivalently replace the constraint with a regularization term. Given a regularization strength $\lambda\geqslant 0$, we can minimize the objective $\mathbb{E}_{1}[\ell_{\alpha}(\kappa\tilde{h}(X),S)]+\lambda(\mathbb{E}_{1}[\tilde{h}(X)^{2}]+\mathbb{E}_{2}[-2\tilde{h}(X)])$ over $\kappa\in\mathbb{R}$ and $\tilde{h}$ square integrable with respect to $\mathbb{P}_{1,X}$. By reparametrizing this optimization problem ($\tilde{h}\mapsto\beta h,\kappa\mapsto 1/\beta$), we may move the scalar parameter $\kappa$ from the first term to the second and third terms. This yields the population-level likelihood ratio regularized quantile regression (LR-QR) objective
$L(h,\beta)=\mathbb{E}_{1}[\ell_{\alpha}(h,S)]+\lambda\mathbb{E}_{1}[\beta^{2}h^{2}]+\lambda\mathbb{E}_{2}[-2\beta h],$ (2)
for some regularization strength $\lambda\geqslant 0$. Finally, since in practice, we cannot optimize over all square-integrable functions, we restrict to functions $h$ in some function class $\mathcal{H}$, leading to the scaled LR-QR problem
$(h^{*},\beta^{*})\in\arg\min_{h\in\mathcal{H},\beta\in\mathbb{R}}L(h,\beta).$ (LR-QR)
We set our threshold function $q$ in (1) to equal $h^{*}$, so that our conformal prediction sets equal $C^{*}(x)=\{y\in\mathcal{Y}:S(x,y)\leqslant h^{*}(x)\}$ for all $x\in\mathcal{X}$.
We solve an empirical version of this objective. We will use our labeled source data $\{(X_{i},Y_{i}):i\in[n_{1}]\}$ to estimate $\mathbb{E}_{1}[\ell_{\alpha}(h,S)]$, our additional unlabeled source data $\mathcal{S}_{3}$ to estimate $\mathbb{E}_{3}[\beta^{2}h^{2}]$, and our unlabeled target data $\mathcal{S}_{2}$ to estimate $\lambda\hat{\mathbb{E}}_{2}[-2\beta h]$. Letting $\hat{\mathbb{E}}_{1}$, $\hat{\mathbb{E}}_{2}$, and $\hat{\mathbb{E}}_{3}$ denote empirical expectations over $\{(X_{i},Y_{i}):i\in[n_{1}]\}$, $\mathcal{S}_{2}$, and $\mathcal{S}_{3}$, respectively, we thus define the empirical risk $\hat{L}(h,\beta)$ as
$\hat{\mathbb{E}}_{1}[\ell_{\alpha}(h,S)]+\lambda\hat{\mathbb{E}}_{3}[\beta^{2}h^{2}]+\lambda\hat{\mathbb{E}}_{2}[-2\beta h],$ (3)
for $\lambda\geqslant 0$. We then solve the following empirical scaled likelihood ratio regularized quantile regression problem:
$(\hat{h},\hat{\beta})\in\arg\min_{h\in\mathcal{H},\beta\in\mathbb{R}}\hat{L}(h,\beta).$ (Empirical-LR-QR)
Our proposed threshold is $q=\hat{h}$. See Algorithm 1.
Algorithm 1 Likelihood-ratio regularized quantile regression
Input: Labeled source data $(x_{i},y_{i})$, $i\in[n_{1}]$, unlabeled target data $\mathcal{S}_{2}$, unlabeled source data $\mathcal{S}_{3}$, Regularization strength: $\lambda\geqslant 0$, Function class: $\mathcal{H}$
1: Compute $S_{i}=S(x_{i},y_{i})$ for all $i\in[n_{1}]$
2: Solve $(\hat{h},\hat{\beta})\leftarrow$ Equation Empirical-LR-QR
Return: $\hat{C}(x)\leftarrow\{y\in\mathcal{Y}:S(x,y)\leqslant\hat{h}(x)\}$
4 Theoretical results
4.1 Infinite sample setting
We first consider the infinite sample or “population” setting, characterizing the solutions of the LR-QR problem from (2) in an idealized scenario where the exact values of the expectations $\mathbb{E}_{1},\mathbb{E}_{2}$ can be calculated. In this case, we will show that if the function class $\mathcal{H}$ is linear and contains the true likelihood ratio $r$, then the optimizer achieves valid coverage in the test domain. Let $r_{\mathcal{H}}$ be the projection of $r$ onto $\mathcal{H}$ in the Hilbert space induced by the inner product $\langle f,g\rangle=\mathbb{E}_{1}[fg]$. The key step is the result below, which characterizes coverage weighted by $r_{\mathcal{H}}$.
Proposition 4.1.
Let $\mathcal{H}$ be a linear function class consisting of square-integrable functions with respect to $\mathbb{P}_{1,X}$. Then under regularity conditions specified in Appendix E (the conditions of Lemma L.3), if $(h^{*},\beta^{*})$ is a minimizer of the objective in Equation LR-QR, then we have
$\mathbb{E}_{1}[r_{\mathcal{H}}(X)\mathbf{1}[S\leqslant h^{*}(X)]]\geqslant 1-\alpha.$
The proof is given in Appendix I. As a consequence of Proposition 4.1, if $\mathcal{H}$ contains the true likelihood ratio $r$, so that $r_{\mathcal{H}}=r$, then in the infinite sample setting, the LR-QR threshold function $h^{*}$ attains valid coverage at test-time:
$\mathbb{E}_{1}[r(X)\mathbf{1}[S\leqslant h^{*}(X)]]=\mathbb{P}_{2}[S\leqslant h^{*}(X)]\geqslant 1-\alpha.$
However, in practice, we can only optimize over finite-dimensional function classes, and as a result we must control the effect of mis-specifying $\mathcal{H}$. If $r$ is not in $\mathcal{H}$, we can derive a lower bound on the coverage as follows. First, write
$\mathbb{E}_{1}[r(X)\mathbf{1}[S(X,Y)\leqslant h^{*}(X)]]$
$=\mathbb{E}_{1}[r_{\mathcal{H}}(X)\mathbf{1}[S(X,Y)\leqslant h^{*}(X)]]+\mathbb{E}_{1}[(r(X)-r_{\mathcal{H}}(X))\mathbf{1}[S(X,Y)\leqslant h^{*}(X)]].$
By Proposition 4.1, the first term on the right-hand side is at least $1-\alpha$. Since the random variable $\mathbf{1}[S(X,Y)\leqslant h^{*}(X)]$ is $\{0,1\}$-valued, the Cauchy-Schwarz inequality implies that the second term on the right-hand side is at least $-\mathbb{E}_{1}[(r-r_{\mathcal{H}})^{2}]^{1/2}$. Thus we have the lower bound
$\mathbb{P}_{2}[Y\in C^{*}(X)]$ $=\mathbb{E}_{1}[r(X)\mathbf{1}[S(X,Y)\leqslant h^{*}(X)]]$
$\geqslant(1-\alpha)-\mathbb{E}_{1}[(r-r_{\mathcal{H}})^{2}]^{1/2}.$
Geometrically, this coverage gap is the result of restricting to $\mathcal{H}$; in fact, $\mathbb{E}_{1}[(r-r_{\mathcal{H}})^{2}]^{1/2}$ is the distance from $r$ to $\mathcal{H}$. This error decreases if $\mathcal{H}$ is made larger, but in the finite sample setting, this comes at the risk of overfitting.
4.2 Finite sample setting
From the analysis of the infinite sample regime, it is clear that if the function class $\mathcal{H}$ is larger, the test-time coverage of the population level LR-QR threshold function $h^{*}$ moves closer to the nominal value. However, in the finite sample setting, optimizing over a larger function class also presents the risk of overfitting. In a finite sample setting, by tuning the regularization parameter $\lambda$, we are trading off the estimation error incurred for the first term of Equation 2, namely $(\hat{\mathbb{E}}_{1}-\mathbb{E}_{1})[\ell_{\alpha}(h(X),S)]$, and the error incurred for the second and third terms of Equation 2, namely $\lambda(\hat{\mathbb{E}}_{3}-\mathbb{E}_{3})[\beta^{2}h^{2}]+\lambda(\hat{\mathbb{E}}_{2}-\mathbb{E}_{2})[-2\beta h]$. Heuristically, for a fixed $h$, the former should be proportional to $1/\sqrt{n_{1}}$, where $n_{1}$ is the number of labeled calibration datapoints, and the latter should be proportional to $\lambda(1/\sqrt{n_{3}}+1/\sqrt{n_{2}})$. Thus, if we pick the regularization to make these two errors of equal order, it will be proportional to $\sqrt{(n_{2}+n_{3})/n_{1}}$.
Put differently, in order to ensure that the Empirical LR-QR threshold $\hat{h}$ from Equation Empirical-LR-QR has valid test coverage, one must choose the regularization $\lambda$ based on the relative amount of labeled and unlabeled data. The unlabeled datapoints carry information about the covariate shift $r$, because $r$ depends only on the distribution of the features. The labeled datapoints provide information about the conditional $(1-\alpha)$-quantile function $q_{1-\alpha}$, which depends only on the conditional distribution of $S|X$. When $\lambda$ is large, our optimization problem places more weight on approximating $r$ (the minimizer of $\mathbb{E}_{1}[(\beta h-r)^{2}]$ in $\beta h$), and if $\lambda$ is small, we instead aim to approximate $q_{1-\alpha}$ (the minimizer of $\ell_{\alpha}(h,S)$ in $h$). Therefore, if the number of unlabeled datapoints ($n_{2}+n_{3}$) is large compared to the number of labeled datapoints ($n_{1}$), our data contains much more information about the covariate shift $r$, and we should set $\lambda$ to be large. If instead $n_{1}$ is very large, the quantile function $q_{1-\alpha}$ can be well-approximated from the labeled calibration datapoints alone, and we set $\lambda$ to be close to zero. In the theoretical results, we make this intuition precise.
In order to facilitate our theoretical analysis in the finite sample setting, we consider constrained versions of Equation LR-QR and Equation Empirical-LR-QR. Fix a collection $\Phi=(\phi_{1},\ldots,\phi_{d})^{\top}$ of $d$ basis functions, where $\phi_{i}:\mathcal{X}\to\mathbb{R}$ for $i\in[d]$. Let $\mathcal{I}=[\beta_{\textnormal{min}},\beta_{\textnormal{max}}]\subset\mathbb{R}$ be a compact interval with $\beta_{\textnormal{min}}>0$. Let $\mathcal{H}_{B}=\{\langle\gamma,\Phi\rangle:\|\gamma\|_{2}\leqslant B<\infty\}$ be the $B$-ball centered at the origin in the linear function class spanned by $\{\phi_{1},\ldots,\phi_{d}\}$. We equip $\mathcal{H}_{B}$ with the norm $\|h\|=\|\gamma\|_{2}$ for $h=\langle\gamma,\Phi\rangle$.
At the population level, consider the following constrained scaled LR-QR problem:
$(h^{*},\beta^{*})\in\arg\min_{h\in\mathcal{H}_{B},\beta\in\mathcal{I}}L(h,\beta).$ (4)
Also consider the following empirical constrained scaled LR-QR problem:
$(\hat{h},\hat{\beta})\in\arg\min_{h\in\mathcal{H}_{B},\beta\in\mathcal{I}}\hat{L}(h,\beta).$ (5)
We begin by bounding the generalization error of an ERM $(\hat{h},\hat{\beta})$ computed via Equation 5.
Theorem 4.2 (Suboptimality gap of ERM for likelihood ratio regularized quantile regression).
Under the regularity conditions specified in Appendix E, and for appropriate choices of the optimization hyperparameters [^2] Specifically, suppose that $\beta_{\textnormal{min}}\leqslant\beta_{\textnormal{lower}}$, $\beta_{\textnormal{max}}\geqslant\beta_{\textnormal{upper}}$, and $B\geqslant B_{\textnormal{upper}}$, where the positive scalars $\beta_{\textnormal{lower}}$, $\beta_{\textnormal{upper}}$, and $B_{\textnormal{upper}}$ are defined in Lemma L.4 in the Appendix, and depend on the data distribution and the choice of basis functions, but not on the data, the sample sizes, or the regularization parameter $\lambda$., for sufficiently large $n_{1},n_{2},n_{3}$, with probability at least $1-\delta$, any optimizer $(\hat{h},\hat{\beta})$ of the empirical constrained LR-QR objective from (5) has suboptimality gap $L(\hat{h},\hat{\beta})-L(h^{*},\beta^{*})$ with respect to the population risk (2) bounded by
$\mathcal{E}_{\textnormal{gen}}:=c\lambda\sqrt{\frac{1}{n_{2}}+\frac{1}{n_{3}}}+\frac{c^{\prime}}{\sqrt{n_{1}}}+\frac{c^{\prime\prime}}{\sqrt{\lambda}}\frac{1}{\sqrt{n_{1}}},$
and $c,c^{\prime},c^{\prime\prime}$ are positive scalars that do not depend on $\lambda$.
The proof is in Appendix J. The generalization error $\mathcal{E}_{\textnormal{gen}}$ is minimized for an optimal regularization on the order of
$\lambda^{*}\propto\frac{1}{n_{1}^{1/3}}\left(\frac{1}{n_{2}}+\frac{1}{n_{3}}\right)^{-1/3},$ (6)
which yields an optimized upper bound of order
$\mathcal{E}_{\textnormal{gen}}^{*}=O\left(\frac{1}{n_{1}^{1/3}}\left(\frac{1}{n_{2}}+\frac{1}{n_{3}}\right)^{1/6}+\frac{1}{\sqrt{n_{1}}}\right).$
As a corollary of Theorem 4.2, we have the following lower bound on the excess marginal coverage of our ERM threshold $\hat{h}$ in the covariate shifted domain. Let $r_{\mathcal{H}_{B}}$ denote the projection of $r$ onto the closed convex set $\mathcal{H}_{B}$ in the Hilbert space induced by the inner product $\langle f,g\rangle=\mathbb{E}_{1}[fg]$.
Theorem 4.3 (Lower bound on marginal coverage of ERM threshold under covariate shift).
Under the same conditions as Theorem 4.2, consider the LR-QR optimizers $\hat{h}$ and $\hat{\beta}$ from (5). Given any $\delta>0$, for sufficiently large $n_{1},n_{2},n_{3}$, we have with probability at least $1-\delta$ that
$\mathbb{P}_{2}\left[Y\in\hat{C}(X)\right]$ $\geqslant(1-\alpha)+2\hat{\beta}\lambda\mathbb{E}_{1}[(r_{\mathcal{H}_{B}}-\hat{\beta}\hat{h})^{2}]-\mathcal{E}_{\textnormal{cov}}-(1-\alpha)\mathbb{E}_{1}[|r-r_{\mathcal{H}_{B}}|],$
where
$\mathcal{E}_{\textnormal{cov}}:=A\left(\frac{1}{n_{2}}+\frac{1}{n_{3}}\right)^{1/4}\lambda+A^{\prime}\frac{\lambda^{-1/4}}{n_{1}^{1/4}}+\Theta\left(\frac{\lambda^{1/2}}{n_{1}^{1/4}}\right),$
and $A,A^{\prime}$ are positive scalars that do not depend on $\lambda$.
The proof is in Appendix K. This result states that our LR-QR method has nearly valid coverage at level $1-\alpha$ under covariate shift, up to small error terms that we can control. The quantity $\mathcal{E}_{\textnormal{cov}}$ vanishes as we collect more data. The term $\mathbb{E}_{1}[|r-r_{\mathcal{H}_{B}}|]$ captures the level of mis-specification by not including the true likelihood ratio function $r$ in our function class $\mathcal{H}_{B}$. This can be decreased by making the function class $\mathcal{H}_{B}$ larger. Of course, this will also increase the size of the terms $A,A^{\prime}$ in our coverage error, but in our theory we show that the dependence is mild. Indeed, the terms depend only on a few geometric properties of $\mathcal{H}_{B}$ such as the eigenvalues of the sample covariance matrix of the basis $\Phi(X)$ under the source distribution, and a quantitative measure of linear dependence of the features; but not explicitly on the dimension of the basis.
We highlight the term $2\hat{\beta}\lambda\mathbb{E}_{1}[(r_{\mathcal{H}_{B}}-\hat{\beta}\hat{h})^{2}]$, which is an error term relating the projected likelihood ratio $r_{\mathcal{H}_{B}}$ to the LR-QR solution $\hat{\beta}\hat{h}$. Crucially, this term is a non-negative quantity multiplied by $\lambda$, and so for appropriate $\lambda$ it may counteract in part the coverage error loss. Consistent with the above observations, we find empirically that choosing small nonzero regularization parameters improves coverage. Moreover, we find that choosing the regularization parameter to be on the order of the optimal value for $\mathcal{E}_{\textnormal{cov}}$ is suitable choice across a range of experiments.
Our proofs are quite involved and require a number of delicate arguments. Crucially, they draw on a novel analysis of coverage via stability bounds from learning theory. Existing stability results cannot directly be applied to our optimization problem, due to our use of a data-dependent regularizer. For instance, in classical settings, the optimal regularization tends to zero as the sample size goes to infinity, but this is not the case for our optimization problem. To overcome this challenge, we combine stability bounds Shalev-Shwartz et al., (2010); Shalev-Shwartz and Ben-David, (2014) with a novel conditioning argument, and we show that the values of $L$ at the minimizers of $\hat{L}$ and $L$ are close by introducing intermediate losses that sequentially swap out empirical expectations $\hat{\mathbb{E}}_{1},\hat{\mathbb{E}}_{2},\hat{\mathbb{E}}_{3}$ with their population counterparts. We then leverage the smoothness of $L$, to derive that the gradient of $L$ at $(\hat{\beta},\hat{h})$ is small. Expanding on our discussion from Section 3.2, we show that a small gradient implies the desired small coverage gap. Our methods may be of independent interest for deriving generalization bounds of empirical risk minimizers with data-dependent regularizers.
5 Experiments
In this section, we compare our method, the LR-QR algorithm, with the following baselines: (1) Split or inductive conformal prediction Papadopoulos et al., (2002); Lei et al., (2018); (2) Weighted conformal prediction Tibshirani et al., (2019), referred to as Weighted-CP; (3) The doubly robust method proposed in Yang et al., (2024), referred to here as 2R-CP; (4) The distributionally robust optimization approach from Cauchois et al., (2024), referred to as DRO-CP.
5.1 Choosing the regularization parameter $\lambda$
Equation 6 suggests an optimal choice of the regularization parameter $\lambda$ in LR-QR. Guided by this theory, we form a grid of ten equally spaced candidate values, ranging from the theoretical $\lambda^{*}$ down to $\frac{\lambda^{*}}{10}$. We then perform three-fold cross-validation over the combined calibration and unlabeled sets (without using any labeled test data) as follows: for each candidate $\lambda$, we train LR-QR with that $\lambda$, and compute a validation measure, namely the $\ell^{2}$-norm of the gradient of the LR-QR objective on the held-out fold. We pick $\lambda$ with the smallest average validation measure across all folds.
This validation measure is motivated by our algorithmic development: the first-order conditions of LR-QR play a fundamental role in ensuring valid coverage in the test domain. While the model is trained to satisfy these conditions on the observed data, we seek to ensure this property generalizes well to unseen data. Thus, our selection criterion is based on two key observations: (1) a small gradient of the LR-QR objective implies reliable coverage, and (2) the regularization parameter $\lambda$ balances the generalization error of the two terms in LR-QR. By minimizing this measure, we select a $\lambda$ that optimally trades off these competing factors.
Finally, we re-train LR-QR on the entire calibration and unlabeled sets using this best $\lambda$, and report coverage and interval size on the held-out labeled test set. This ensures that no test labels are used during hyper-parameter tuning.
Additionally, in Appendix B, we provide deeper insight on different regimes of regularization in practice through an ablation study.
5.2 Communities and Crime
We evaluate our methods on the Communities and Crime dataset Redmond, (2002), which contains $1994$ datapoints corresponding to communities in the United States, along with socio-economic and demographic statistics. The task is to predict the per-capita violent crime rate (a real-valued response) from a 127-dimensional input.
Figure 1: (Left) Coverage. (Right) Average prediction set size on the Communities and Crime dataset.
Setup. We first randomly select half of the data as a training set, and use it to fit a ridge regression model $\hat{f}(\cdot)$ as our predictor. We tune the ridge regularizer with five-fold cross-validation. We use the remaining half to design four covariate shift scenarios, determined by the frequency of a specific racial subgroup (Black, White, Hispanic, and Asian). For each of these features, we find the median value $m$ over the remaining dataset. Datapoints with feature value at most $m$ form our source set, while the rest form our target set. This creates a covariate shift between calibration and test, as the split procedure only observes the covariates, and is independent of labels. We then further split the target set into roughly equal unlabeled and labeled subsets. The unlabeled subset and the calibration data (only their covariates, not the labels) is used for estimating $r$, whereas the labeled subset of test is held out only for final evaluation. The same procedure is applied to each of the four racial subgroups, creating four distinct partitions.
Conformity score and likelihood ratio estimation. The conformity score is the absolute residual $(x,y)\mapsto s(x,y)\;=\;|y-\hat{f}(x)\,|.$ Several baselines require an estimate of the likelihood ratio $r$, which we obtain by training a logistic regression model to distinguish datapoints in the unlabeled target data from those in the source data. We then use $\hat{r}=\hat{p}/\bigl{(}1-\hat{p}\bigr{)}$, where $x\mapsto\hat{p}(x)$ is the predicted probability that $x$ belongs to the target distribution.
Experimental details. We set the function class $\mathcal{H}$ to be linear maps from the feature space to $\mathbb{R}$. The regularization $\lambda$ is tuned as explained in Section 5.1. All experimental results are averaged over 1000 random splits.
Results. Figure 1 depicts the performance of all methods. Notably, split conformal undercovers in two setups and overcovers in the other two. Methods relying on estimating the likelihood ratio and DRO fail to track the nominal coverage precisely, particularly in the first setup on the left. However, LR-QR is closer to the nominal level of coverage, showing a stronger adaptivity to the covariate shift.
5.3 RxRx1 data - WILDS
Figure 2: (Left) Coverage, (Right) Average prediction set size on the RxRx1 dataset from the WILDS repository.
Our next experiment uses the RxRx1 dataset Sypetkowski et al., (2023) from the WILDS repository Koh et al., (2021), which is designed to evaluate model robustness under distribution shifts. The RxRx1 task involves classifying cell images based on 1,339 laboratory genetic treatments. These images, captured using fluorescent microscopy, originate from 51 independent experiments. Variations in execution and environmental conditions lead to systematic differences across experiments, affecting the distribution of input features (e.g., lighting, cell morphology) while the relationship between inputs and labels remains unchanged. This situation creates covariate shift where the marginal distribution of inputs shifts across domains, but the conditional distribution $\mathbb{P}_{Y|X}$ remains the same.
For this task, we use a ResNet50 model He et al., (2016) trained by the WILDS authors on 37 of the 51 experiments. Using data from the other 14 experiments, we construct 14 distinct evaluations. In each evaluation, one experiment is selected as the target dataset, and its data is evenly split into an unlabeled target set and a labeled test set. The labeled data from the other 13 experiments serves as the source dataset. We evaluate all baselines across these 14 settings.
Conformity score and likelihood ratio estimation. The nonconformity score $S(x,y)$ is the negative log-likelihood $(x,y)\mapsto-\log f_{x}(y),$ where $f_{x}(y)$ is the probability the pretrained model assigns to the image-label pair $(x,y)$. Several baselines require an estimate of $r$, which we obtain by training a logistic regression model on top of the representation layer of the pretrained model to distinguish datapoints in the unlabeled target data from those in the source data. We then use $\hat{r}=\hat{p}/\bigl{(}1-\hat{p}\bigr{)}$, where $x\mapsto\hat{p}(x)$ is the predicted probability that $x$ came from the target distribution.
Experimental details. For our proposed method, we set the function class $\mathcal{H}$ to be a linear head on top of the representation layer of the pretrained model. The $\lambda$ is tuned using the cross-validation explained in Section 5.1. Also, all experimental results are averaged over 50 random splits.
Results. Figure 2 presents the coverage and average prediction set size for all methods. To enhance visual interpretability, we display results for eight randomly selected settings out of the 14, with the full plot provided in Figure 3 in the Appendix. The x-axis shows the indices of the test experimental condition. As shown, LR-QR consistently adheres more closely to the nominal coverage value of $0.9$ compared to other baselines. The average absolute deviations of coverage from the nominal value in terms of percentage for LR-QR, 2R-CP, weighted-CP, DRO-CP, and Split conformal are $1.56,2.62,4.03,5.25$, and $5.36$, respectively.
Notably, split conformal prediction, which assumes exchangeability between calibration and test data, shows under- and overcoverage due to the shift. The coverage of weighted CP and 2R-CP is also far from the target, showing that directly estimating the likelihood ratio and conditional quantile is insufficient to correct the coverage violations in the case of high-dimensional image data. Further, the superior coverage of LR-QR is not due to inflated prediction sets, as the sizes are on par with other methods.
6 Discussion and future work
Distribution shifts are inevitable in machine learning applications. Consequently, precise uncertainty quantification under distribution shifts is essential to ensuring the safety and reliability of predictive models in practice. This challenge becomes even more pronounced when dealing with high-dimensional data, where classical statistical procedures often fail to generalize effectively. In this work, we develop a new conformal prediction method, which we call LR-QR, designed to provide valid test-time coverage under covariate shifts between calibration and test data. In contrast to existing approaches in the literature, LR-QR avoids directly estimating the likelihood ratio function between calibration and test time. Instead, it leverages certain one-dimensional projections of the likelihood ratio function, which effectively enhance LR-QR’s performance in high-dimensional tasks compared to other baselines.
While this paper primarily focuses on marginal test-time coverage guarantees, we acknowledge that in many practical scenarios, marginal guarantees alone may not suffice. An interesting direction for future work is to explore whether the techniques and intuitions developed here can be extended to provide stronger conditional guarantees at test time in the presence of covariate shifts. In particular, is it possible to achieve group-conditional coverage at test time (e.g., see Bastani et al., (2022); Jung et al., (2023); Gibbs et al., (2023)) without directly estimating the likelihood ratio function?
Additionally, several open questions remain regarding the regularization technique in LR-QR. Specifically, what alternative forms of regularization, beyond the mean squared error used in this work, could be employed to further improve test-time coverage? Which type of regularization is optimal in the sense that it yields the most precise test-time coverage? Furthermore, what is the most effective strategy for tuning the regularization strength? In particular, can these ideas be extended to design a hyperparameter-free algorithm? Finally, the data-adaptive regularization introduced in this work may have applications beyond conformal prediction, serving as a general technique to improve robustness to covariate shifts in other machine learning problems.
7 Acknowledgments
ED and SJ were supported by NSF, ARO, ONR, AFOSR, and the Sloan Foundation. The work of HH, SK, and GP was supported by the NSF Institute for CORE Emerging Methods in Data Science (EnCORE).
Appendix A Additional figures
Figure 3: (Left) Coverage, (Right) Average prediction set size.
Figure 4: Ablation study on the effect of $\lambda$ on LR-QR performance in the experimental setup of Section 5.2. In (a), the theoretically suggested regime for $\lambda$ effectively ensures valid test-time coverage. Additionally, in (b), the average norm of the gradients reaches its lowest value in the regime predicted by theory, highlighting the effectiveness of the cross-validation procedure described in Section 5.1.
Appendix B Ablation studies
Here we provide an ablation study for $\lambda$, the regularization strength that appears in the LR-QR objective. In the same regression setup as Section 5.2, instead of selecting $\lambda$ via cross-validation, here we sweep the value of $\lambda$ from 00 to $2$, and we plot the coverage of the LR-QR algorithm on the test data. Here, note that the split ratios between train, calibration, and test (both labeled and unlabeled data) are fixed and similar to the setup in Section 5.2. We report the averaged plots over 100 independent splits.
Figure 4(a) displays the effect of different regimes of $\lambda$. At one extreme, when $\lambda$ is close to zero, the LR-QR algorithm reduces to ordinary quantile regression. In this regime, the LR-QR algorithm behaves similarly to the algorithm from Gibbs et al., (2023), without the test covariate imputation. In other words, when we set $\lambda=0$, we try to provide coverage with respect to all the covariate shifts in the linear function class that we optimize over. As we can see in Figure 4(a), this can lead to overfitting and undercoverage of the test labels. As we increase $\lambda$, as a direct effect of the regularization, the coverage gap decreases. This is primarily due to the fact that larger $\lambda$ restricts the space of quantile regression optimization in such a way that it does not hurt the test time coverage, since the regularization is designed to shrink the optimization space towards the true likelihood-ratio. Thus, the regularization improves the generalization of the selected threshold, as the effective complexity of the function class is getting smaller. That being said, this phenomenon is only applicable if $\lambda$ lies within a certain range; once $\lambda$ grows too large, due to the data-dependent nature of our regularization, the generalization error of the regularization term itself becomes non-negligible and hinders the precise test-time coverage of the LR-QR threshold. As is highlighted in Figure 4(a), our theoretical results suggest an optimal regime for $\lambda$ which can best exploit the geometric properties of the LR-QR threshold.
Additionally, Figure 4(b) demonstrates the effectiveness of the cross-validation technique described in Section 5.1. We sweep the value of $\lambda$ from 00 to $2$ and plot the average norm of the gradient on the holdout sets for the cross-validation procedure explained in Section 5.1. As our theory suggests, it is now evident that the stationary conditions of LR-QR are closely tied to the valid test-time coverage of our method. For all values of $\lambda$, during training, we fit the LR-QR objective to the data, ensuring that the average norm of the gradients is zero. However, when evaluating the LR-QR objective on the holdout set, the average norm of the gradients is no longer zero due to generalization errors. Selecting $\lambda$ correctly minimizes this generalization error, thereby providing more precise test-time coverage.
Appendix C Further related work
The basic concept of prediction sets dates back to foundational works such as Wilks, (1941), Wald, (1943), Scheffe and Tukey, (1945), and Tukey, (1947, 1948). The early ideas of conformal prediction were developed in Saunders et al., (1999); Vovk et al., (1999). With the rise of machine learning, conformal prediction has emerged as a widely used framework for constructing prediction sets (e.g., Papadopoulos et al., , 2002; Vovk et al., , 2005; Lei et al., , 2018; Angelopoulos and Bates, , 2021). Since then, efforts have been emerged to improve prediction set size efficiency (e.g., Sadinle et al., , 2019; Stutz et al., , 2022; Bai et al., , 2022; Kiyani et al., 2024b, ; Noorani et al., , 2024) and conditional coverage guarantees (e.g., Foygel Barber et al., , 2021; Sesia and Romano, , 2021; Gibbs et al., , 2023; Romano et al., , 2019; Kiyani et al., 2024a, ; Jung et al., , 2023)
Numerous works have addressed conformal prediction under various types of distribution shift (Tibshirani et al., , 2019; Park et al., 2022b, ; Park et al., 2022a, ; Qiu et al., , 2023; Si et al., , 2024). For example, Tibshirani et al., (2019) and Lei and Candès, (2021), investigated conformal prediction under covariate shift, assuming the likelihood ratio between source and target covariates is known or can be precisely estimated from data. Park et al., 2022b developed prediction sets with a calibration-set conditional (PAC) property under covariate shift. Qiu et al., (2023); Yang et al., (2024) developed prediction sets with asymptotic coverage that are doubly robust in the sense that their coverage error is bounded by the product of the estimation errors of the quantile function of the score and the likelihood ratio. Cauchois et al., (2024) construct prediction sets based on a distributionally robust optimization approach. Gui et al., (2024) develop methods based on an isotonic regression estimate of the likelihood ratio. Qin et al., (2024) combine a parametric working model with a resampling approach to construct prediction sets under covariate shift. Bhattacharyya and Barber, (2024) analyze weighted conformal prediction in the special case of covariate shifts defined by a finite number of groups. Ai and Ren, (2024) reweight samples to adapt to covariate shift, while simultaneously using distributionally robust optimization to protect against worst-case joint distribution shifts. Kasa et al., (2024) construct prediction sets by using unlabeled test data to modify the score function used for conformal prediction.
Appendix D Notation and conventions
Constants are allowed to depend on dimension only through properties of the population and sample covariance matrices of the features, and the amount of linear independence of the features; see the quantities $\lambda_{\textnormal{min}}(\Sigma)$, $\lambda_{\textnormal{max}}$, $c_{\textnormal{min}}$, $c_{\textnormal{max}}$, and $c_{\textnormal{indep}}$ defined in Appendix E. In Landau notation ($o$, $O$, $\Theta$), we hide constants. We say that a sequence of events holds with high probability if the probability of the events tends to unity. We define $\mathcal{S}_{1}$ as the features of the labeled calibration dataset. All functions that we minimize can readily be verified to be continuous, and thus attain a minimum over the compact domains over which we minimize them; thus all our minimizers will be well-defined. We may not mention this further. We denote by $\mathbf{1}[A]$ the indicator of an event $A$. Recall that $\mathcal{H}$ denotes the linear function class =$\mathcal{H}=\{\langle\gamma,\Phi\rangle:\gamma\in\mathbb{R}^{d}\}$. This defines a one-to-one correspondence between $\mathbb{R}^{d}$ and $\mathcal{H}$. This enables us to view functions defined on $\mathbb{R}^{d}$ equivalently as defined on $\mathcal{H}$. In our analysis, we will use such steps without further discussion. Unless stated otherwise, $\mathcal{H}$ is equipped with the norm $\|h\|:=\|\gamma\|_{2}$ for $h=\langle\gamma,\Phi\rangle$. Given a differentiable function $\varphi:\mathcal{H}\to\mathbb{R}$, its directional derivative at $f=\langle\gamma,\Phi\rangle\in\mathcal{H}$ in the direction defined by the function $g\in\mathcal{H}$ is defined as $\frac{d}{d\varepsilon}\big{|}_{\varepsilon=0}\varphi(f+\varepsilon g)$. Note that if we write $g=\langle\tilde{\gamma},\Phi\rangle$ for some $\tilde{\gamma}\in\mathbb{R}^{d}$, then the directional derivative of $\varphi$ at $f$ equals $\langle\tilde{\gamma},\nabla_{\gamma}\varphi(\gamma)\rangle$, where $\nabla_{\gamma}\varphi(\gamma)$ denotes the gradient of $\varphi:\mathbb{R}^{d}\to\mathbb{R}$ evaluated at $\gamma\in\mathbb{R}^{d}$.
Appendix E Conditions
Condition 1.
Suppose $C_{\Phi}=\sup_{x\in\mathcal{X}}\|\Phi(x)\|_{2}$ is finite.
Condition 2.
For the population covariance matrix $\Sigma=\mathbb{E}_{1}[\Phi\Phi^{\top}]$, we have $\lambda_{\textnormal{min}}(\Sigma)>0$ and $\lambda_{\textnormal{max}}(\Sigma)$ is of constant order, not depending on the sample size, or any other problem parameter.
Condition 3.
For the sample covariance matrix $\hat{\Sigma}=\frac{1}{n_{3}}\sum_{k=1}^{n_{3}}\Phi(x_{k})\Phi(x_{k})^{\top}$, we have both $\lambda_{\textnormal{min}}(\hat{\Sigma})\geqslant c_{\textnormal{min}}>0$ and $\lambda_{\textnormal{max}}(\hat{\Sigma})\leqslant c_{\textnormal{max}}$ of constant order with probability $1-o(n_{3}^{-1})$.
Condition 4.
Defining $C_{1}$ as in (10) in Appendix H, assume there exists an upper bound $C_{1,\textnormal{upper}}$ on $\mathbb{E}[C_{1}]$ of constant order.
Condition 5.
The conditional density $f_{S|X=x}$ exists for all $x\in\mathcal{X}$, and $C_{f}=\sup_{x\in\mathcal{X}}\|f_{S|X=x}(s)\|_{\infty}$ is a finite constant.
The following can be interpreted as an independence assumption on the basis functions.
Condition 6.
Suppose $\inf_{v\in S^{d-1}}\mathbb{E}_{1}[|\langle v,\Phi\rangle|]\geqslant c_{\textnormal{indep}}>0$ for some constant $c_{\textnormal{indep}}$.
Condition 7.
Suppose $\frac{\mathbb{E}_{1}[rh_{0}^{*}]}{\mathbb{E}_{1}[|h_{0}^{*}|^{2}]^{1/2}}\geqslant c_{\textnormal{align}}>0$ for some minimizer $h_{0}^{*}$ of the objective in Equation 21 with regularization $\lambda=0$.
Condition 8.
Suppose $\mathbb{E}_{1}[r^{2}]$ is finite.
Condition 9.
The constant function $h:\mathcal{X}\to\mathbb{R}$ given by $h(x)=1$ for all $x\in\mathcal{X}$ is in $\mathcal{H}$.
The following ensures that the zero function $0\in\mathcal{H}$ is not a minimizer of the objective in Equation LR-QR.
Condition 10.
For each $\lambda\geqslant 0$, there exists $h\in\mathcal{H}$ and $\beta\in\mathbb{R}$ such that
$\mathbb{E}_{1}[\ell_{\alpha}(h,S)]+\lambda\mathbb{E}_{1}[(\beta h-r)^{2}]<\mathbb{E}_{1}[\ell_{\alpha}(0,S)]+\lambda\mathbb{E}_{1}[r^{2}].$
Appendix F Constants
The following are the constants that appear in Theorem 4.2:
$\rho_{1}:=2\beta_{\textnormal{max}}^{2}BC_{\Phi}^{2}+2\beta_{\textnormal{max}}C_{\Phi},\quad\mu_{1}:=2\beta_{\textnormal{min}}^{2}c_{\textnormal{min}},\quad\rho_{2}:=(1-\alpha)C_{\Phi},$
$\widetilde{C}_{1}:=\frac{4\rho_{1}^{2}}{\mu_{1}},\quad\widehat{C}_{2}:=\frac{4\rho_{2}^{2}}{2\beta_{\textnormal{min}}^{2}c_{\textnormal{min}}},\quad A_{1}:=\sqrt{\frac{64\widetilde{C}_{1}a_{1}}{\delta}},\quad A_{2}:=\sqrt{\frac{128\widehat{C}_{2}a_{2}}{\delta}}.$
Further,
$A_{3}$ $:=(1-\alpha)(BC_{\Phi}+1)\sqrt{\frac{1}{2}\log\frac{8}{\delta}},\qquad A_{4}:=\sqrt{2}(\beta_{\textnormal{max}}BC_{\Phi})\sqrt{\frac{1}{2}\log\frac{16}{\delta}}\max\left\{\beta_{\textnormal{max}}BC_{\Phi},4\right\},$
$A_{5}$ $:=A_{1}+A_{4},\qquad a_{1}:=2C_{\Phi}(C_{2,\textnormal{upper}}+C_{2,\textnormal{max}})(1+\beta_{\textnormal{max}}BC_{\Phi}),\qquad a_{2}:=(1-\alpha)C_{\Phi}(C_{1,\textnormal{upper}}+C_{1,\textnormal{max}}).$
The following are the constants that appear in Theorem 4.3:
$A_{6}:=2\beta_{\textnormal{max}}^{2}\sqrt{4B^{2}\lambda_{\textnormal{max}}(\Sigma)},\qquad A_{7}:=\sqrt{4B^{2}\beta_{\textnormal{max}}^{2}\lambda_{\textnormal{max}}(\Sigma)},\qquad A_{8}:=\sqrt{2B^{2}C_{f}\lambda_{\textnormal{max}}(\Sigma)},\qquad A_{9}:=A_{6}+A_{7},$
and
$A_{10}$ $:=A_{9}A_{5}^{1/2},\qquad A_{11}:=\max\{A_{9}A_{3}^{1/2},A_{8}A_{5}^{1/2}\},\qquad A_{12}:=A_{9}A_{2}^{1/2},\qquad A_{13}:=A_{8}A_{3}^{1/2},\qquad A_{14}:=A_{8}A_{2}^{1/2}.$
Appendix G Generalization bound for regularized loss
The following is a generalization of Shalev-Shwartz and Ben-David, (2014, Corollary 13.6).
Lemma G.1 (Generalization bound for regularized loss; extension of Shalev-Shwartz and Ben-David, (2014)).
Fix a compact and convex function class $\tilde{\mathcal{H}}$ equipped with a norm $\|\cdot\|_{\tilde{\mathcal{H}}}$, a compact interval $\mathcal{I}\subseteq\mathbb{R}$, and a sample space $\mathcal{Z}$. Consider the objective function $f:\tilde{\mathcal{H}}\times\mathcal{I}\times\mathcal{Z}\to\mathbb{R}$ given by $(h,\beta,z)\mapsto f(h,\beta,z):=\mathcal{J}(h,\beta,z)+\mathcal{R}(h,\beta)$, where $\mathcal{R}:\tilde{\mathcal{H}}\times\mathcal{I}\to\mathbb{R}$ is a regularization function, and $\mathcal{J}:\tilde{\mathcal{H}}\times\mathcal{I}\times\mathcal{Z}\to\mathbb{R}$ can be decomposed as $\mathcal{J}(h,\beta,z):=\mathcal{J}_{1}(h,\beta,z_{1})+\mathcal{J}_{2}(h,\beta,z_{2})$ for two functions $\mathcal{J}_{1},\mathcal{J}_{2}:\tilde{\mathcal{H}}\times\mathcal{I}\times\mathcal{Z}\to\mathbb{R}$.
Given distributions $\mathcal{D}_{1},\mathcal{D}_{2}$ on $\mathcal{Z}$, let $\mathcal{L}:\tilde{\mathcal{H}}\times\mathcal{I}\to\mathbb{R}$ be given for all $h,\beta$ by
$\mathcal{L}(h,\beta)=\mathbb{E}_{Z_{1}\sim\mathcal{D}_{1},Z_{2}\sim\mathcal{D}_{2}}[f(h,\beta,Z_{1},Z_{2})]$
denote the population risk, averaging over independent datapoints $Z_{1}\sim\mathcal{D}_{1}$ and $Z_{2}\sim\mathcal{D}_{2}$. Suppose that for both $Z\sim\mathcal{D}_{1}$ and $Z\sim\mathcal{D}_{2}$, $|\mathcal{J}_{1}(h,\beta,Z)|$ and $|\mathcal{J}_{2}(h,\beta,Z)|$ are almost surely bounded by a quantity not depending on $h\in\tilde{\mathcal{H}}$ and $\beta\in\mathcal{I}$.
Let $\hat{\mathcal{L}}:\tilde{\mathcal{H}}\times\mathcal{I}\to\mathbb{R}$ denote the empirical risk computed over $Z_{i,1}\overset{i.i.d.}{\sim}\mathcal{D}_{1}$, $i\in[m_{1}]$ and $Z_{j,2}\overset{i.i.d.}{\sim}\mathcal{D}_{2}$, $j\in[m_{2}]$, given by
^$\hat{\mathcal{L}}(h,\beta):=\frac{1}{m_{1}}\sum_{i=1}^{m_{1}}\mathcal{J}_{1}(h,\beta,Z_{i,1})+\frac{1}{m_{2}}\sum_{j=1}^{m_{2}}\mathcal{J}_{2}(h,\beta,Z_{j,2})+\mathcal{R}(h,\beta).$
Assume that for each fixed $\beta\in\mathcal{I}$ and $z\in\mathcal{Z}$,
  • $h\mapsto\mathcal{J}_{1}(h,\beta,z)$ is convex and $\rho$-Lipschitz with respect to the norm $\|\cdot\|_{\tilde{\mathcal{H}}}$,
  • $h\mapsto\mathcal{J}_{2}(h,\beta,z)$ is convex and $\rho$-Lipschitz with respect to the norm $\|\cdot\|_{\tilde{\mathcal{H}}}$, and
  • $h\mapsto\hat{\mathcal{L}}(h,\beta)$ is $\mu$-strongly convex with respect to the norm $\|\cdot\|_{\tilde{\mathcal{H}}}$ with probability $1-o(m_{1}^{-1}+m_{2}^{-1})$,
where the deterministic values $\mu=\mu(\beta)$ and $\rho=\rho(\beta)$ may depend on $\beta$.
Let $(\hat{h},\hat{\beta})$ denote an ERM, i.e., a minimizer of ^$\hat{\mathcal{L}}(h,\beta)$ over $\tilde{\mathcal{H}}\times\mathcal{I}$. Let $\hat{h}_{\beta}$ denote a minimizer of the empirical risk in $h$ for fixed $\beta$.
Suppose the stochastic process $\beta\mapsto W_{\beta}$ given by $W_{\beta}=\mathcal{L}(\hat{h}_{\beta},\beta)-\hat{\mathcal{L}}(\hat{h}_{\beta},\beta)$ for $\beta\in\mathcal{I}$ obeys $|W_{\beta}-W_{\beta^{\prime}}|\leqslant K|\beta-\beta^{\prime}|$ for all $\beta,\beta^{\prime}\in\mathcal{I}$ for some random variable $K$, and suppose that the probability of $K_{m_{1},m_{2}}\leqslant K_{\textnormal{max}}$ converges to unity as $m_{1},m_{2}\to\infty$, for some constant $K_{\textnormal{max}}$. Suppose that there exists a constant $C>0$ such that for all $\beta\in\mathcal{I}$,
$\frac{4\rho(\beta)^{2}}{\mu(\beta)}\leqslant C.$ (7)
Then for sufficiently large $m_{1},m_{2}$, with probability at least $1-\delta$,
$|\mathcal{L}(\hat{h},\hat{\beta})-\hat{\mathcal{L}}(\hat{h},\hat{\beta})|\leqslant\sqrt{\frac{16CK_{\textnormal{max}}}{\delta}(m_{1}^{-1}+m_{2}^{-1})}.$
Remark 1.
A special case is when we do not have any data from $\mathcal{D}_{2}$, and instead all $m_{1}$ datapoints are sampled i.i.d. from $\mathcal{D}_{1}$. In this case, defining with a slight abuse of notation $\mathcal{J}:=\mathcal{J}_{1}$, the statement simplifies to the analysis of the empirical risk
^$\hat{\mathcal{L}}(h,\beta):=\frac{1}{m_{1}}\sum_{i=1}^{m_{1}}\mathcal{J}(h,\beta,Z_{i,1})+\mathcal{R}(h,\beta).$
If for each fixed $\beta\in\mathcal{I}$, we have that $h\mapsto\mathcal{J}(h,\beta,z)$ is convex and $\rho$-Lipschitz with respect to the norm $\|\cdot\|_{\tilde{\mathcal{H}}}$, and if $|\mathcal{J}(h,\beta,Z)|$ is almost surely bounded by a quantity not depending on $h\in\tilde{\mathcal{H}}$ and $\beta\in\mathcal{I}$ for $Z\sim\mathcal{D}_{1}=\mathcal{D}_{2}$, then under the remaining assumptions, we obtain the slightly stronger bound
$|\mathcal{L}(\hat{h},\hat{\beta})-\hat{\mathcal{L}}(\hat{h},\hat{\beta})|\leqslant\sqrt{\frac{16CK_{\textnormal{max}}}{\delta m_{1}}}.$
We omit the proof, because it is exactly as below.
Remark 2.
We relax the strong convexity assumption on the regularizer $\mathcal{R}$ from Shalev-Shwartz and Ben-David, (2014, Corollary 13.6), substituting it with the less restrictive condition of strong convexity of the empirical loss $\hat{\mathcal{L}}$. In order to use assumptions that merely hold with high probability, we impose a boundedness condition on $\mathcal{J}$.
Proof.
Fix $\beta$ and let $E$ denote the event that $h\mapsto\hat{\mathcal{L}}(h,\beta)$ is $\mu$-strongly convex in $h$. By assumption, $E$ occurs with probability $1-o(m_{1}^{-1}+m_{2}^{-1})$.
We modify the proof of Shalev-Shwartz and Ben-David, (2014, Corollary 13.6) as follows. Let $Z_{1}^{\prime}\sim\mathcal{D}_{1}$ and $Z_{2}^{\prime}\sim\mathcal{D}_{2}$ be drawn independently from all other randomness. For a fixed $i\in[m_{1}]$, let $h\mapsto\hat{\mathcal{L}}_{i,1}(h,\beta)$ denote the empirical risk computed from the sample $(Z_{1,1},\ldots,Z_{i-1,1},Z_{1}^{\prime},Z_{i+1,1},\ldots,Z_{m_{1},1})\cup(Z_{1,2},\ldots,Z_{m_{2},2})$, and let $\hat{h}_{\beta}^{(i)}$ denote an ERM for this sample. Let $I$ be drawn from $[m_{1}]$ uniformly at random. The variables $J,\hat{\mathcal{L}}_{J,2}(h,\beta),\hat{h}_{\beta}^{(J)}$ are defined similarly but for the sample from $\mathcal{D}_{2}$.
Note that for fixed $\beta$, similarly to the argument in Shalev-Shwartz and Ben-David, (2014, Theorem 13.2), we have
$\mathbb{E}[\mathcal{L}(\hat{h}_{\beta},\beta)]$ $=\mathbb{E}_{Z_{1}^{\prime}\sim\mathcal{D}_{1},Z_{2}^{\prime}\sim\mathcal{D}_{2}}[\mathcal{J}_{1}(\hat{h}_{\beta},\beta,Z_{1}^{\prime})+\mathcal{J}_{2}(\hat{h}_{\beta},\beta,Z_{2}^{\prime})+\mathcal{R}(\hat{h}_{\beta},\beta)]$
$=\mathbb{E}_{Z_{1}^{\prime}\sim\mathcal{D}_{1},Z_{2}^{\prime}\sim\mathcal{D}_{2}}[\mathcal{J}_{1}(\hat{h}_{\beta}^{(I)},\beta,Z_{I,1})+\mathcal{J}_{2}(\hat{h}_{\beta}^{(J)},\beta,Z_{J,2})+\mathcal{R}(\hat{h}_{\beta},\beta)]$
and
$\mathbb{E}[\hat{\mathcal{L}}(\hat{h}_{\beta},\beta)]=\mathbb{E}[\mathcal{J}_{1}(\hat{h}_{\beta},\beta,Z_{I,1})+\mathcal{J}_{2}(\hat{h}_{\beta},\beta,Z_{J,2})+\mathcal{R}(\hat{h}_{\beta},\beta)].$
Therefore
$\mathbb{E}[\mathcal{L}(\hat{h}_{\beta},\beta)-\hat{\mathcal{L}}(\hat{h}_{\beta},\beta)]=(\mathbb{E}[\mathcal{J}_{1}(\hat{h}_{\beta}^{(I)},\beta,Z_{I,1})-\mathcal{J}_{1}(\hat{h}_{\beta},\beta,Z_{I,1})])+(\mathbb{E}[\mathcal{J}_{2}(\hat{h}_{\beta}^{(J)},\beta,Z_{J,2})-\mathcal{J}_{2}(\hat{h}_{\beta},\beta,Z_{J,2})]).$
Further, splitting the expectations over $E$ and its complement $E^{c}$, this further equals
$(\mathbb{E}[(\mathcal{J}_{1}(\hat{h}_{\beta}^{(I)},\beta,Z_{I,1})-\mathcal{J}_{1}(\hat{h}_{\beta},\beta,Z_{I,1}))\cdot\mathbf{1}[E]]+\mathbb{E}[(\mathcal{J}_{1}(\hat{h}_{\beta}^{(I)},\beta,Z_{I,1})-\mathcal{J}_{1}(\hat{h}_{\beta},\beta,Z_{I,1}))\cdot\mathbf{1}[E^{c}]])$ (8)
$+(\mathbb{E}[(\mathcal{J}_{2}(\hat{h}_{\beta}^{(J)},\beta,Z_{J,2})-\mathcal{J}_{2}(\hat{h}_{\beta},\beta,Z_{J,2}))\cdot\mathbf{1}[E]]+\mathbb{E}[(\mathcal{J}_{2}(\hat{h}_{\beta}^{(J)},\beta,Z_{J,2})-\mathcal{J}_{2}(\hat{h}_{\beta},\beta,Z_{J,2}))\cdot\mathbf{1}[E^{c}]]).$
On the event $E$, $h\mapsto\hat{\mathcal{L}}(h,\beta)$ is $\mu$-strongly convex. Now, consider the setting of Shalev-Shwartz and Ben-David, (2014, Corollary 13.6). We claim that the arguments in their proof hold if we replace the regularizer $h\mapsto\lambda\|h\|^{2}$ by $h\mapsto\mathcal{R}(h,\beta)$, as they only leverage the strong convexity of the overall empirical loss $\hat{\mathcal{L}}$. Indeed, working on the event $E$, since $\hat{\mathcal{L}}$ is $\mu$-strongly convex, we have that ^$\hat{\mathcal{L}}(h)-\hat{\mathcal{L}}(\hat{h}_{\beta})\geqslant\frac{1}{2}\mu\|h-\hat{h}_{\beta}\|^{2}$ for all $h\in\tilde{\mathcal{H}}$. Next, for any $h_{1},h_{2}\in\tilde{\mathcal{H}}$, we have
^$\hat{\mathcal{L}}(h_{2})-\hat{\mathcal{L}}(h_{1})=\hat{\mathcal{L}}_{I,1}(h_{2})-\hat{\mathcal{L}}_{I,1}(h_{1})+\frac{\mathcal{J}_{1}(h_{2},\beta,Z_{I,1})-\mathcal{J}_{1}(h_{1},\beta,Z_{I,1})}{m_{1}}-\frac{\mathcal{J}_{1}(h_{2},\beta,Z_{1}^{\prime})-\mathcal{J}_{1}(h_{1},\beta,Z_{1}^{\prime})}{m_{1}}.$
Setting $h_{2}=\hat{h}_{\beta}^{(I)}$ and $h_{1}=\hat{h}$, since $\hat{h}_{\beta}^{(I)}$ minimizes $h\mapsto\hat{\mathcal{L}}_{I,1}(h,\beta)$, and using our lower bound on ^$\hat{\mathcal{L}}(h)-\hat{\mathcal{L}}(\hat{h}_{\beta})$, we deduce
$\frac{1}{2}\mu\|\hat{h}_{\beta}^{(I)}-\hat{h}_{\beta}\|^{2}\leqslant\frac{\mathcal{J}_{1}(\hat{h}_{\beta}^{(I)},\beta,Z_{I,1})-\mathcal{J}_{1}(\hat{h}_{\beta},\beta,Z_{I,1})}{m_{1}}-\frac{\mathcal{J}_{1}(\hat{h}_{\beta}^{(I)},\beta,Z_{1}^{\prime})-\mathcal{J}_{1}(\hat{h}_{\beta},\beta,Z_{1}^{\prime})}{m_{1}}.$ (9)
Since by assumption, $h\mapsto\mathcal{J}_{1}(h,\beta,z)$ is $\rho$-Lipschitz, we have the bounds $|\mathcal{J}_{1}(\hat{h}_{\beta}^{(I)},\beta,Z_{I,1})-\mathcal{J}_{1}(\hat{h}_{\beta},\beta,Z_{I,1})|\leqslant\rho|\hat{h}_{\beta}^{(I)}-\hat{h}_{\beta}|$ and $|\mathcal{J}_{1}(\hat{h}_{\beta}^{(I)},\beta,Z_{1}^{\prime})-\mathcal{J}_{1}(\hat{h}_{\beta},\beta,Z_{1}^{\prime})|\leqslant\rho|\hat{h}_{\beta}^{(I)}-\hat{h}_{\beta}|$. Plugging these into Equation 9, we obtain $\frac{1}{2}\mu\|\hat{h}_{\beta}^{(I)}-\hat{h}_{\beta}\|^{2}\leqslant\frac{2\rho}{m_{1}}\|\hat{h}_{\beta}^{(I)}-\hat{h}_{\beta}\|$, so that $\|\hat{h}_{\beta}^{(I)}-\hat{h}_{\beta}\|\leqslant\frac{4\rho(\beta)}{\mu(\beta)m_{1}}$. Using once again that $h\mapsto\mathcal{J}_{1}(h,\beta,z)$ is $\rho$-Lipschitz, we find $|\mathcal{J}_{1}(\hat{h}_{\beta}^{(I)},\beta,Z_{I,1})-\mathcal{J}_{1}(\hat{h}_{\beta},\beta,Z_{I,1})|\leqslant\frac{4\rho(\beta)^{2}}{\mu(\beta)m_{1}}$.
Similarly, on the event $E$, we have the bound $|\mathcal{J}_{2}(\hat{h}_{\beta}^{(J)},\beta,Z_{J,2})-\mathcal{J}_{2}(\hat{h}_{\beta},\beta,Z_{J,2})|\leqslant\frac{4\rho(\beta)^{2}}{\mu(\beta)m_{2}}$. Thus the first and third terms are bounded in magnitude by $\frac{4\rho(\beta)^{2}}{\mu(\beta)m_{1}}$ and $\frac{4\rho(\beta)^{2}}{\mu(\beta)m_{2}}$, respectively. Due to (7), their sum is at most $C(m_{1}^{-1}+m_{2}^{-1})$.
By our assumption that $|\mathcal{J}_{1}(h,\beta,Z)|$ and $|\mathcal{J}_{2}(h,\beta,Z)|$ are almost surely bounded by a constant for both $Z\sim\mathcal{D}_{1}$ and $Z\sim\mathcal{D}_{2}$, and our assumption that $\mathbb{P}\left[E^{c}\right]=o(m_{1}^{-1}+m_{2}^{-1})$, the second term and fourth terms from (8) sum to $o(m_{1}^{-1}+m_{2}^{-1})$. Thus for for each $\beta$, for sufficiently large $m_{1},m_{2}$, we have $\mathbb{E}[|W_{\beta}|]\leqslant 2C(m_{1}^{-1}+m_{2}^{-1})$. By Markov’s inequality, for any fixed $t>0$, $|W_{\beta}|>t$ with probability at most $\frac{2C}{t}(m_{1}^{-1}+m_{2}^{-1})$. We now use chaining. Let $N$ be an $\varepsilon$-net for $\mathcal{I}$. Then using the fact that by assumption, the process $W$ is $K_{m_{1},m_{2}}$-Lipschitz, and by a union bound,
$\mathbb{P}\left[\sup_{\beta\in\mathcal{I}}|W_{\beta}|>K_{m_{1},m_{2}}\varepsilon+t\right]\leqslant\mathbb{P}\left[\sup_{\beta\in N}|W_{\beta}|>t\right]\leqslant|N|\frac{2C}{t}(m_{1}^{-1}+m_{2}^{-1}).$
Pick $N$ with $|N|=1/\varepsilon$, and set $t=\frac{4C}{\delta}(m_{1}^{-1}+m_{2}^{-1})\frac{1}{\varepsilon}$. We deduce that
$\sup_{\beta\in\mathcal{I}}|W_{\beta}|>K_{m_{1},m_{2}}\varepsilon+\frac{4C}{\delta}(m_{1}^{-1}+m_{2}^{-1})\frac{1}{\varepsilon}$
with probability at most $\frac{\delta}{2}$. Set $\varepsilon=\sqrt{\frac{4C}{K_{m_{1},m_{2}}\delta}(m_{1}^{-1}+m_{2}^{-1})}$. We deduce that
$\sup_{\beta\in\mathcal{I}}|W_{\beta}|>\sqrt{\frac{16CK_{m_{1},m_{2}}}{\delta}(m_{1}^{-1}+m_{2}^{-1})}$
with probability at most $\frac{\delta}{2}$. Since the probability of $K_{m_{1},m_{2}}\leqslant K_{\textnormal{max}}$ converges to unity, for sufficiently large $m_{1},m_{2}$,
$\sup_{\beta\in\mathcal{I}}|W_{\beta}|>\sqrt{\frac{16CK_{\textnormal{max}}}{\delta}(m_{1}^{-1}+m_{2}^{-1})}$
holds with probability at most $\delta$. Since $|W_{\hat{\beta}}|\leqslant\sup_{\beta\in\mathcal{I}}|W_{\beta}|$, we may conclude. ∎
Appendix H Lipschitz process
Lemma H.1 (Lipschitzness of minimizer of perturbed strongly convex objective).
Let $\mathcal{C}\subseteq\mathbb{R}^{d}$ be a closed convex set. Suppose $\psi:\mathcal{C}\to\mathbb{R}$ is $\mu$-strongly convex and $g:\mathcal{C}\to\mathbb{R}$ is $L$-smooth. Suppose also that $\psi+g$ is convex. Let $x_{\psi}$ denote the minimizer of $\psi$ in $\mathcal{C}$, and let $x_{\psi+g}$ denote the minimizer of $\psi+g$ in $\mathcal{C}$. Then for any $x\in\mathcal{C}$,
$\|x_{\psi+g}-x_{\psi}\|_{2}\leqslant\frac{1}{\mu}(L\|x_{\psi+g}-x\|_{2}+\|\nabla g(x)\|_{2}).$
Proof.
Since $\psi$ is $\mu$-strongly convex and since $x_{\psi+g},x_{\psi}$ are minimizers of $\psi+g,\psi$ respectively,
$\mu\|x_{\psi+g}-x_{\psi}\|_{2}^{2}$ $\leqslant\langle\nabla\psi(x_{\psi+g})-\nabla\psi(x_{\psi}),x_{\psi+g}-x_{\psi}\rangle$
$=\langle\nabla(\psi+g)(x_{\psi+g}),x_{\psi+g}-x_{\psi}\rangle+\langle\nabla\psi(x_{\psi}),x_{\psi}-x_{\psi+g}\rangle-\langle\nabla g(x_{\psi+g}),x_{\psi+g}-x_{\psi}\rangle$
$\leqslant-\langle\nabla g(x_{\psi+g}),x_{\psi+g}-x_{\psi}\rangle$
$=-\langle\nabla g(x_{\psi+g})-\nabla g(x),x_{\psi+g}-x_{\psi}\rangle-\langle\nabla g(x),x_{\psi+g}-x_{\psi}\rangle,$
so that by $L$-smoothness of $g$,
$\mu\|x_{\psi+g}-x_{\psi}\|_{2}^{2}\leqslant(L\|x_{\psi+g}-x\|_{2}+\|\nabla g(x)\|_{2})\|x_{\psi+g}-x_{\psi}\|_{2},$
which implies the result. ∎
Lemma H.2 (Lipschitzness of minimizer of perturbed ERM).
[…]
$C_{1}$ $=\left(\beta_{\textnormal{min}}^{2}\lambda_{\textnormal{min}}(\hat{\Sigma}))^{-1}((2\beta_{\textnormal{max}}\lambda_{\textnormal{max}}(\hat{\Sigma})B+C_{\Phi})+4\beta_{\textnormal{max}}\lambda_{\textnormal{max}}(\hat{\Sigma})B\right),\qquad C_{2}=B+\beta_{\textnormal{max}}C_{1}.$ (10)
Proof.
First, consider $\hat{h}_{\beta}$. Fix $\beta>\beta^{\prime}$ in $\mathcal{I}$. Recalling the definition of $\hat{L}$ from (3), the difference between the objectives $\hat{L}(h,\beta)$ and $\hat{L}(h,\beta^{\prime})$ is the quadratic
$g(h):=\hat{L}(h,\beta)-\hat{L}(h,\beta^{\prime})=\lambda\hat{\mathbb{E}}_{3}[(\beta^{2}-(\beta^{\prime})^{2})h^{2}]+\lambda\hat{\mathbb{E}}_{2}[-2(\beta-\beta^{\prime})h].$
We claim that $g$ is $2\lambda(\beta^{2}-(\beta^{\prime})^{2})\lambda_{\textnormal{min}}(\hat{\Sigma})$-strongly convex and $2\lambda(\beta^{2}-(\beta^{\prime})^{2})\lambda_{\textnormal{max}}(\hat{\Sigma})$-smooth in $h$. To see this, write $h=\langle\gamma,\Phi\rangle$ for $\gamma\in\mathbb{R}^{d}$, and note that $g$ can be rewritten as
$g(\gamma)=\lambda(\beta^{2}-(\beta^{\prime})^{2})\gamma^{\top}\hat{\Sigma}\gamma-2(\beta-\beta^{\prime})\lambda\gamma^{\top}\hat{\mathbb{E}}_{2}[\Phi],$
a quadratic whose Hessian equals $2\lambda(\beta^{2}-(\beta^{\prime})^{2})\hat{\Sigma}$, which implies the claim.
Similarly, we claim that the function $\psi(h):=\hat{L}(h,\beta^{\prime})$ is $2\lambda(\beta^{\prime})^{2}\lambda_{\textnormal{min}}(\hat{\Sigma})$-strongly convex in $h$. To see this, again write $h=\langle\gamma,\Phi\rangle$ for $\gamma\in\mathbb{R}^{d}$, and note that $\psi$ can be rewritten as
$\psi(\gamma)=\lambda(\beta^{\prime})^{2}\gamma^{\top}\hat{\Sigma}\gamma+\hat{\mathbb{E}}_{1}[\ell_{\alpha}(\gamma^{\top}\Phi,S)]+\lambda\hat{\mathbb{E}}_{2}[-2\beta\gamma^{\top}\Phi].$
By Lemma N.3, the second term is convex, and since the third term is linear, it too is convex. The Hessian of the quadratic first term is $2\lambda(\beta^{\prime})^{2}\hat{\Sigma}$, from which it follows that $\psi$ is $2\lambda(\beta^{\prime})^{2}\lambda_{\textnormal{min}}(\hat{\Sigma})$-strongly convex.
Thus $\psi$ and $g$ satisfy the conditions of Lemma H.1, which implies the bound
$\|\hat{h}_{\beta}-\hat{h}_{\beta^{\prime}}\|\leqslant(2\lambda(\beta^{\prime})^{2}\lambda_{\textnormal{min}}(\hat{\Sigma}))^{-1}(\|\nabla g(\hat{h}_{g})\|_{2}+2\lambda(\beta^{2}-(\beta^{\prime})^{2})\lambda_{\textnormal{max}}(\hat{\Sigma})\cdot\|\hat{h}_{\beta}-\hat{h}_{g}\|),$ (11)
where $\hat{h}_{g}=\hat{h}_{g,\beta,\beta^{\prime}}$ denotes the minimizer of $g$ in $\mathcal{H}_{B}$. Since
$\nabla g(\gamma)=\lambda(\beta-\beta^{\prime})((\beta+\beta^{\prime})2\hat{\Sigma}\gamma-2\hat{\mathbb{E}}_{2}[\Phi]),$
and by $|\beta|,|\beta^{\prime}|\leqslant\beta_{\textnormal{max}}$, $\|\gamma\|\leqslant B$, and 1, we have
$\|\nabla g(\gamma)\|_{2}\leqslant\lambda(4\beta_{\textnormal{max}}\lambda_{\textnormal{max}}(\hat{\Sigma})B+2C_{\Phi})|\beta-\beta^{\prime}|$
for $\beta,\beta^{\prime}\in\mathcal{I}$ and $h\in\mathcal{H}_{B}$. Plugging this into the bound (11) on $\|\hat{h}_{\beta}-\hat{h}_{\beta^{\prime}}\|$ and using the fact that $\beta^{\prime}\geqslant\beta_{\textnormal{min}}$ and $\|\hat{h}_{\beta}\|,\|\hat{h}_{g}\|\leqslant B$,
$\|\hat{h}_{\beta}-\hat{h}_{\beta^{\prime}}\|\leqslant(2\lambda\beta_{\textnormal{min}}^{2}\lambda_{\textnormal{min}}(\hat{\Sigma}))^{-1}\left(\lambda(4\beta_{\textnormal{max}}\lambda_{\textnormal{max}}(\hat{\Sigma})B+2C_{\Phi})|\beta-\beta^{\prime}|+8\lambda\beta_{\textnormal{max}}\lambda_{\textnormal{max}}(\hat{\Sigma})B|\beta-\beta^{\prime}|\right)$
Thus we may take
$C_{1}=(\beta_{\textnormal{min}}^{2}\lambda_{\textnormal{min}}(\hat{\Sigma}))^{-1}\left((2\beta_{\textnormal{max}}\lambda_{\textnormal{max}}(\hat{\Sigma})B+C_{\Phi})+4\beta_{\textnormal{max}}\lambda_{\textnormal{max}}(\hat{\Sigma})B\right).$
For the map $\beta\mapsto\beta\hat{h}_{\beta}$, fix $\beta>\beta^{\prime}$ in $\mathcal{I}$, and write $\|\beta\hat{h}_{\beta}-\beta^{\prime}\hat{h}_{\beta^{\prime}}\|\leqslant|\beta-\beta^{\prime}|\|\hat{h}_{\beta}\|+|\beta^{\prime}|\|\hat{h}_{\beta}-\hat{h}_{\beta^{\prime}}\|$. For the first term, note that since $\hat{h}_{\beta}\in\mathcal{H}_{B}$ implies $\|\hat{h}_{\beta}\|\leqslant B$, the first term is bounded by $B|\beta-\beta^{\prime}|$. For the second term, note that since $|\beta^{\prime}|\leqslant\beta_{\textnormal{max}}$ and since $\beta\mapsto\hat{h}_{\beta}$ is $C_{1}$-Lipschitz on $\mathcal{I}$, the second term is bounded by $\beta_{\textnormal{max}}C_{1}|\beta-\beta^{\prime}|$. Summing, we deduce that $\beta\mapsto\beta\hat{h}_{\beta}$ is $C_{2}$-Lipschitz on $\mathcal{I}$, where $C_{2}=B+\beta_{\textnormal{max}}C_{1}$. ∎
Lemma H.3 (Lipschitzness of minimizer of perturbed auxiliary ERM).
Under 1, we have that $\beta\mapsto\tilde{h}_{\beta}$ is $C_{1}$-Lipschitz on $\mathcal{I}$, and $\beta\mapsto\beta\tilde{h}_{\beta}$ is $C_{2}$-Lipschitz on $\mathcal{I}$.
Proof.
The proof is almost identical to Lemma H.2. ∎
Recalling $c_{\textnormal{min}}$ and $c_{\textnormal{max}}$ from 3, define
$C_{1,\textnormal{max}}$ $=(\beta_{\textnormal{min}}^{2}c_{\textnormal{min}})^{-1}((2\beta_{\textnormal{max}}c_{\textnormal{max}}B+C_{\Phi})+4\beta_{\textnormal{max}}c_{\textnormal{max}}B),\qquad C_{2,\textnormal{max}}=B+\beta_{\textnormal{max}}C_{1,\textnormal{max}},$ (12)
so that by 3, $C_{1}\leqslant C_{1,\textnormal{max}}$ and $C_{2}\leqslant C_{2,\textnormal{max}}$ with probability tending to unity over the randomness in $\mathcal{S}_{3}$.
We now compute the Lipschitz constants of the processes used in the proof of Theorem 4.2.
Recall $\bar{L}$ from (16), $\hat{L}$ from (3), $\tilde{L}$ from (15), and $\mathcal{H}_{B}=\{\langle\gamma,\Phi\rangle:\|\gamma\|_{2}\leqslant B<\infty\}$ from Section 4. For any fixed $\beta\in\mathcal{I},$ define $\hat{h}_{\beta}$ as the minimizer of $h\mapsto\hat{L}(h,\beta)$ over $\mathcal{H}_{B}$, which exists under the conditions of Theorem 4.2 due to our argument checking the convexity of $h\mapsto\hat{L}(h,\beta)$ in Term (I) in the proof of Theorem 4.2.
Lemma H.4.
Assume the conditions of Theorem 4.2. Define the stochastic processes $\bar{W}_{\beta}$ and $\tilde{W}_{\beta}$ on $\mathcal{I}$ given by $\beta\mapsto(\bar{L}-\hat{L})(\hat{h}_{\beta},\beta)$ and $\beta\mapsto(\tilde{L}-\hat{L})(\hat{h}_{\beta},\beta)$, respectively. Then $\bar{W}_{\beta}$ is $K_{1,\lambda}$-Lipschitz on $\mathcal{I}$ with probability tending to unity as $n_{1},n_{2},n_{3}\to\infty$, and $\tilde{W}_{\beta}$ is $K_{2,\lambda}$-Lipschitz on $\mathcal{I}$ with probability tending to unity as $n_{1},n_{2},n_{3}\to\infty$, where
$K_{1,\lambda}:=2C_{\Phi}(C_{2,\textnormal{upper}}+C_{2,\textnormal{max}})(1+\beta_{\textnormal{max}}BC_{\Phi})\lambda=:a_{1}\lambda,\qquad K_{2,\lambda}:=(1-\alpha)C_{\Phi}(C_{1,\textnormal{upper}}+C_{1,\textnormal{max}})=:a_{2}.$
with $C_{1,\textnormal{max}}$ and $C_{2,\textnormal{max}}$ are defined in (12) and where $C_{1,\textnormal{upper}}$ satisfies 4 and $C_{2,\textnormal{upper}}:=BC_{\Phi}+\beta_{\textnormal{max}}C_{\Phi}C_{1,\textnormal{upper}}$. In fact, $\bar{W}$ is $K_{1,\lambda}$-Lipschitz on $\mathcal{I}$ with probability tending to unity conditional on $\mathcal{S}_{1}$, and $\tilde{W}$ is $K_{2,\lambda}$-Lipschitz on $\mathcal{I}$ deterministically, when conditioning on $\mathcal{S}_{2},\mathcal{S}_{3}$, when the event $C_{1}\leqslant C_{1,\textnormal{max}}$ holds.
Proof.
We start with the process $\tilde{W}$. Consider $\beta,\beta^{\prime}\in\mathcal{I}$. Note that for any $(h,\beta)$, using the definition of $\tilde{L}$ from (15), we have the identity
$\tilde{L}(h,\beta)-\hat{L}(h,\beta)=\mathbb{E}_{1}[\ell_{\alpha}(h,S)]-\hat{\mathbb{E}}_{1}[\ell_{\alpha}(h,S)].$
Thus we may write
$\tilde{W}_{\beta}-\tilde{W}_{\beta^{\prime}}$ $=(\mathbb{E}_{1}[\ell_{\alpha}(\hat{h}_{\beta},S)]-\mathbb{E}_{1}[\ell_{\alpha}(\hat{h}_{\beta^{\prime}},S)])-(\hat{\mathbb{E}}_{1}[\ell_{\alpha}(\hat{h}_{\beta},S)]-\hat{\mathbb{E}}_{1}[\ell_{\alpha}(\hat{h}_{\beta^{\prime}},S)]),$
so that
$|\tilde{W}_{\beta}-\tilde{W}_{\beta^{\prime}}|$ $\leqslant\mathbb{E}_{1}[|\ell_{\alpha}(\hat{h}_{\beta},S)-\ell_{\alpha}(\hat{h}_{\beta^{\prime}},S)|]+\hat{\mathbb{E}}_{1}[|\ell_{\alpha}(\hat{h}_{\beta},S)-\ell_{\alpha}(\hat{h}_{\beta^{\prime}},S)|]$ (13)
Note that we have the uniform bound
$|\ell_{\alpha}(\hat{h}_{\beta},S)-\ell_{\alpha}(\hat{h}_{\beta^{\prime}},S)|\leqslant(1-\alpha)|\hat{h}_{\beta}-\hat{h}_{\beta^{\prime}}|\leqslant(1-\alpha)C_{\Phi}\|\hat{h}_{\beta}-\hat{h}_{\beta^{\prime}}\|\leqslant(1-\alpha)C_{\Phi}C_{1}|\beta-\beta^{\prime}|,$
where in the first step we applied Lemma N.2, in the second step we used 1 to apply Lemma N.4, and in the third step we used Lemma H.2. Thus the first term in Equation 13 is bounded by $(1-\alpha)C_{\Phi}\mathbb{E}_{1}[C_{1}]|\beta-\beta^{\prime}|$, and the second term in Equation 13 is bounded by $(1-\alpha)C_{\Phi}\hat{\mathbb{E}}_{1}[C_{1}]|\beta-\beta^{\prime}|$. Summing, we deduce that
$|\tilde{W}_{\beta}-\tilde{W}_{\beta^{\prime}}|\leqslant(1-\alpha)C_{\Phi}(\mathbb{E}_{1}[C_{1}]+\hat{\mathbb{E}}_{1}[C_{1}])|\beta-\beta^{\prime}|,$
so that the process $\tilde{W}$ is $K_{2}$-Lipschitz with $K_{2}:=(1-\alpha)C_{\Phi}(\mathbb{E}_{1}[C_{1}]+\hat{\mathbb{E}}_{1}[C_{1}])$.
We now condition on $\mathcal{S}_{2},\mathcal{S}_{3}$. Observe that $C_{1},C_{2}$ are $\mathcal{S}_{3}$-measurable (as $\hat{\Sigma}$ from Condition 3 is $\mathcal{S}_{3}$-measurable). Since $\mathbb{E}_{1}[C_{1}]\leqslant C_{1,\textnormal{upper}}$, on the event that $C_{1}\leqslant C_{1,\textnormal{max}}$, we have $K_{2}\leqslant K_{2,\lambda}$, where $K_{2,\lambda}=(1-\alpha)C_{\Phi}(C_{1,\textnormal{upper}}+C_{1,\textnormal{max}})$, as claimed.
We now continue with the process $\bar{W}$. Consider $\beta,\beta^{\prime}\in\mathcal{I}$. Note that for any $(h,\beta)$, using the definition of $\bar{L}$ from Equation 16, we have the identity
$\bar{L}(h,\beta)-\hat{L}(h,\beta)=(\lambda\mathbb{E}_{3}[\beta^{2}h^{2}]+\lambda\mathbb{E}_{2}[-2\beta h])-(\lambda\hat{\mathbb{E}}_{3}[\beta^{2}h^{2}]+\lambda\hat{\mathbb{E}}_{2}[-2\beta h]).$
Thus we may write
$\bar{W}_{\beta}-\bar{W}_{\beta^{\prime}}$ $=\lambda(\mathbb{E}_{3}[\beta^{2}\hat{h}_{\beta}^{2}]-\mathbb{E}_{3}[(\beta^{\prime})^{2}\hat{h}_{\beta^{\prime}}^{2}])+\lambda(\mathbb{E}_{2}[-2\beta\hat{h}_{\beta}]-\mathbb{E}_{2}[-2\beta^{\prime}\hat{h}_{\beta^{\prime}}])$
$-\lambda(\hat{\mathbb{E}}_{3}[\beta^{2}\hat{h}_{\beta}^{2}]-\hat{\mathbb{E}}_{3}[(\beta^{\prime})^{2}\hat{h}_{\beta^{\prime}}^{2}])-\lambda(\hat{\mathbb{E}}_{2}[-2\beta\hat{h}_{\beta}]-\hat{\mathbb{E}}_{2}[-2\beta^{\prime}\hat{h}_{\beta^{\prime}}]),$
so that
$|\bar{W}_{\beta}-\bar{W}_{\beta^{\prime}}|$ $\leqslant\lambda\mathbb{E}_{3}[|\beta^{2}\hat{h}_{\beta}^{2}-(\beta^{\prime})^{2}\hat{h}_{\beta^{\prime}}^{2}|]+2\lambda\mathbb{E}_{2}[|\beta\hat{h}_{\beta}-\beta^{\prime}\hat{h}_{\beta^{\prime}}|]$
$+\lambda\hat{\mathbb{E}}_{3}[|\beta^{2}\hat{h}_{\beta}^{2}-(\beta^{\prime})^{2}\hat{h}_{\beta^{\prime}}^{2}|]+2\lambda\hat{\mathbb{E}}_{2}[|\beta\hat{h}_{\beta}-\beta^{\prime}\hat{h}_{\beta^{\prime}}|]$ (14)
The integrands of the first and third terms of Appendix H can be uniformly bounded as
$|\beta^{2}\hat{h}_{\beta}^{2}-(\beta^{\prime})^{2}\hat{h}_{\beta^{\prime}}^{2}|\leqslant|\beta\hat{h}_{\beta}-\beta^{\prime}\hat{h}_{\beta^{\prime}}|\cdot|\beta\hat{h}_{\beta}+\beta^{\prime}\hat{h}_{\beta^{\prime}}|\leqslant C_{\Phi}\|\beta\hat{h}_{\beta}-\beta^{\prime}\hat{h}_{\beta^{\prime}}\|\cdot C_{\Phi}\|\beta\hat{h}_{\beta}+\beta^{\prime}\hat{h}_{\beta^{\prime}}\|$
$\leqslant C_{\Phi}C_{2}|\beta-\beta^{\prime}|\cdot 2C_{\Phi}\beta_{\textnormal{max}}B=2\beta_{\textnormal{max}}BC_{\Phi}^{2}C_{2}|\beta-\beta^{\prime}|.$
where in the first step we used difference of squares, in the second step we used 1 to apply Lemma N.4, in the third step we applied Lemma H.2 to bound the first factor and the triangle inequality and the bounds $\beta\leqslant\beta_{\textnormal{max}}$ for $\beta\in\mathcal{I}$ and $\|h\|\leqslant B$ for $h\in\mathcal{H}_{B}$ to bound the second factor. The integrand of the second and fourth term in (H) can be bounded as $|\beta\hat{h}_{\beta}-\beta^{\prime}\hat{h}_{\beta^{\prime}}|\leqslant C_{\Phi}\|\beta\hat{h}_{\beta}-\beta^{\prime}\hat{h}_{\beta^{\prime}}\|\leqslant C_{\Phi}C_{2}|\beta-\beta^{\prime}|$, where in the first step we used 1 to apply Lemma N.4, and in the second step we applied Lemma H.2.
Plugging these into our bound in Appendix H, we deduce
$|\bar{W}_{\beta}-\bar{W}_{\beta^{\prime}}|\leqslant(2C_{\Phi}(\mathbb{E}_{2}[C_{2}]+\hat{\mathbb{E}}_{2}[C_{2}])+2\beta_{\textnormal{max}}C_{\Phi}^{2}B(\mathbb{E}_{3}[C_{2}]+\hat{\mathbb{E}}_{3}[C_{2}]))\lambda|\beta-\beta^{\prime}|,$
so that the process $\bar{W}$ is $K_{1}$-Lipschitz with
$K_{1}$ $=(2C_{\Phi}(\mathbb{E}_{2}[C_{2}]+\hat{\mathbb{E}}_{2}[C_{2}])+2\beta_{\textnormal{max}}C_{\Phi}^{2}B(\mathbb{E}_{3}[C_{2}]+\hat{\mathbb{E}}_{3}[C_{2}]))\lambda.$
We now work conditional on $\mathcal{S}_{1}$. On the event that $C_{1}\leqslant C_{1,\textnormal{max}}$ and $C_{2}\leqslant C_{2,\textnormal{max}}$, and by 4, we have $K_{1}\leqslant K_{1,\textnormal{max}}$, where
$K_{1,\lambda}$ $=(2C_{\Phi}(C_{2,\textnormal{upper}}+C_{2,\textnormal{max}})+2\beta_{\textnormal{max}}C_{\Phi}^{2}B(C_{2,\textnormal{upper}}+C_{2,\textnormal{max}}))\lambda$
$=2C_{\Phi}(C_{2,\textnormal{upper}}+C_{2,\textnormal{max}})(1+\beta_{\textnormal{max}}BC_{\Phi})\lambda.$
Since $C_{1}\leqslant C_{1,\textnormal{max}}$ and $C_{2}\leqslant C_{2,\textnormal{max}}$ with probability tending to one due to Condition 3, $K_{1}\leqslant K_{1,\lambda}$ and $K_{2}\leqslant K_{2,\lambda}$ both hold with probability tending to one if we uncondition on $\mathcal{S}_{1}$, and we are done. ∎
Appendix I Proof of Proposition 4.1
Fix $\lambda\geqslant 0$. Under the assumptions of Lemma L.3, there exists a global minimizer $(h^{*},\beta^{*})$ of $L(h,\beta)$. The first order condition with respect to $\beta$ reads $02\lambda\mathbb{E}_{1}[h^{*}(X)(\beta^{*}h^{*}(X)-r(X))]=$ 0. By Lemma N.5, the first order condition with respect to $h$ reads
$\mathbb{E}_{1}[h^{*}(X)(\mathbb{P}_{S|X}[S(X,Y)\leqslant h^{*}(X)]-(1-\alpha))]+2\lambda\mathbb{E}_{1}[\beta^{*}h(X)(\beta^{*}h^{*}(X)-r(X))]=0$
for all $h\in\mathcal{H}$. Setting $h=r_{\mathcal{H}}$ in the second equation, and subtracting $(\beta^{*})^{2}$ times the first equation from the second, we deduce that
$\mathbb{E}_{1}[h^{*}(X)(\mathbb{P}_{S|X}[S(X,Y)\leqslant h^{*}(X)]-(1-\alpha))]$
$+2\lambda\mathbb{E}_{1}[\beta^{*}\cdot r_{\mathcal{H}}(X)\cdot(\beta^{*}h^{*}(X)-r(X))]-2\lambda\mathbb{E}_{1}[\beta^{*}\cdot\beta^{*}h^{*}(X)\cdot(\beta^{*}h^{*}(X)-r(X))]$
$=\mathbb{E}_{1}[h^{*}(X)(\mathbb{P}_{S|X}[S(X,Y)\leqslant h^{*}(X)]-(1-\alpha))]+2\lambda\mathbb{E}_{1}[\beta^{*}(r_{\mathcal{H}}(X)-\beta^{*}h^{*}(X))(\beta^{*}h^{*}(X)-r(X))]$
$=\mathbb{E}_{1}[h^{*}(X)\mathbb{P}_{S|X}[S(X,Y)\leqslant h^{*}(X)]]-(1-\alpha)-2\lambda\beta^{*}\mathbb{E}_{1}[(r_{\mathcal{H}}(X)-\beta^{*}h^{*}(X))^{2}]=0.$
Therefore,
$\mathbb{E}_{1}[r_{\mathcal{H}}(X)\mathbb{P}_{S|X}[S(X,Y)\leqslant h^{*}(X)]]=(1-\alpha)+2\lambda\beta^{*}\mathbb{E}_{1}[(r_{\mathcal{H}}(X)-\beta^{*}h^{*}(X))^{2}],$
which implies the result.
Appendix J Proof of Theorem 4.2
Recall that $\mathcal{S}_{1}$ are the features of the labeled calibration dataset. We also recall the notation $\mathbb{E}_{j}$ and $\hat{\mathbb{E}}_{j}$ for $j=1,2,3$ from Section 2. Given the unlabeled test data $\mathcal{S}_{2}$ and the unlabeled calibration data $\mathcal{S}_{3}$, define the auxiliary risks for $h\in\mathcal{H}_{B},\beta\in\mathcal{I}$,
$\tilde{L}(h,\beta;\mathcal{S}_{2},\mathcal{S}_{3}):=\mathbb{E}_{1}[\ell_{\alpha}(h,S)]+\lambda\hat{\mathbb{E}}_{3}[\beta^{2}h^{2}]+\lambda\hat{\mathbb{E}}_{2}[-2\beta h]$ (15)
and
$\bar{L}(h,\beta;\mathcal{S}_{1}):=\hat{\mathbb{E}}_{1}[\ell_{\alpha}(h,S)]+\lambda\mathbb{E}_{3}[\beta^{2}h^{2}]+\lambda\mathbb{E}_{2}[-2\beta h].$ (16)
Let
$(\tilde{h},\tilde{\beta})\in\arg\min_{h\in\mathcal{H}_{B},\beta\in\mathcal{I}}\tilde{L}(h,\beta;\mathcal{S}_{2},\mathcal{S}_{3}).$ (17)
For convenience, we leave implicit the dependence of $\tilde{L}$ and $(\tilde{h},\tilde{\beta})$ on $\mathcal{S}_{2}$, $\mathcal{S}_{3}$ and the dependence of $\bar{L}$ on $\mathcal{S}_{1}$.
In order to study the generalization error, we write
$L(\hat{h},\hat{\beta})-L(h^{*},\beta^{*})$ $=(L(\hat{h},\hat{\beta})-\tilde{L}(\hat{h},\hat{\beta}))+(\tilde{L}(\hat{h},\hat{\beta})-\hat{L}(\hat{h},\hat{\beta}))+(\hat{L}(\hat{h},\hat{\beta})-\hat{L}(\tilde{h},\tilde{\beta}))$
$+(\hat{L}(\tilde{h},\tilde{\beta})-\tilde{L}(\tilde{h},\tilde{\beta}))+(\tilde{L}(\tilde{h},\tilde{\beta})-\tilde{L}(h^{*},\beta^{*}))+(\tilde{L}(h^{*},\beta^{*})-L(h^{*},\beta^{*})).$
Since $(\hat{h},\hat{\beta})$ is a minimizer of the risk $\hat{L}$, we have $\hat{L}(\hat{h},\hat{\beta})-\hat{L}(\tilde{h},\tilde{\beta})\leqslant 0$, and since $(\tilde{h},\tilde{\beta})$ is a minimizer of the risk $\tilde{L}$, we have $\tilde{L}(\tilde{h},\tilde{\beta})-\tilde{L}(h^{*},\beta^{*})\leqslant 0$. Thus our generalization error is bounded by the remaining four terms:
$L(\hat{h},\hat{\beta})-L(h^{*},\beta^{*})$ $\leqslant(L(\hat{h},\hat{\beta})-\tilde{L}(\hat{h},\hat{\beta}))+(\tilde{L}(\hat{h},\hat{\beta})-\hat{L}(\hat{h},\hat{\beta}))$
$+(\hat{L}(\tilde{h},\tilde{\beta})-\tilde{L}(\tilde{h},\tilde{\beta}))+(\tilde{L}(h^{*},\beta^{*})-L(h^{*},\beta^{*}))$
$=:(I)+(II)+(III)+(IV).$ (18)
We study the generalization error by conditioning on the unlabeled calibration or test data. Then our regularization becomes data-independent. Conditional on $\mathcal{S}_{1}$, Term (I) can be handled with Lemma G.1 above. Conditional on $\mathcal{S}_{2},\mathcal{S}_{3}$, Term (II) can be handled with Lemma G.1 above. Terms (III) and (IV) are empirical processes at fixed functions, conditional on $\mathcal{S}_{2},\mathcal{S}_{3}$.
Term (I): We work conditional on $\mathcal{S}_{1}$. First, note that due to the definition of $\hat{L}$ from (3), we can write for any $(h,\beta)$,
$L(h,\beta)-\tilde{L}(h,\beta)=\bar{L}(h,\beta)-\hat{L}(h,\beta).$
Since $\bar{L}(h,\beta)-\hat{L}(h,\beta)$ can be viewed as a difference of a population risk $\lambda\mathbb{E}_{3}[\beta^{2}h^{2}]+\lambda\mathbb{E}_{2}[-2\beta h]$ and an empirical risk $\lambda\hat{\mathbb{E}}_{3}[\beta^{2}h^{2}]+\lambda\hat{\mathbb{E}}_{2}[-2\beta h]$ with “regularizer” $\hat{\mathbb{E}}_{1}[\ell_{\alpha}(h,S)]$, this expression enables us to apply Lemma G.1 to bound $\bar{L}(\hat{h},\hat{\beta})-\hat{L}(\hat{h},\hat{\beta})$.
Explicitly, we can write
$\frac{1}{\lambda}\hat{L}(h,\beta)=\hat{\mathbb{E}}_{3}[\beta^{2}h^{2}]+\hat{\mathbb{E}}_{2}[-2\beta h]+\frac{1}{\lambda}\hat{\mathbb{E}}_{1}[\ell_{\alpha}(h,S)].$
Hence, fixing $\beta$, we can apply Lemma G.1, choosing $m_{1}=n_{3}$ and $m_{2}=n_{2}$. Further, we choose ~$\tilde{\mathcal{H}}:=\mathcal{H}_{B}=\{\langle\gamma,\Phi\rangle:\|\gamma\|_{2}\leqslant B<\infty\}$ with the norm $\langle\gamma,\Phi\rangle=\|\gamma\|_{2}$. Moreover, letting $z=(x^{\prime\prime},x^{\prime})$ for $x^{\prime\prime},x^{\prime}\in\mathcal{X}$, and $\xi=1/\lambda$, we use the objective function given by $(h,z)\mapsto f_{1}(h,z)=\mathcal{J}(h,\beta,z)+\mathcal{R}(h,\beta)$, where $\mathcal{J}(h,\beta,z)=\mathcal{J}_{1}(h,\beta,z)+\mathcal{J}_{2}(h,\beta,z)$, and where
$\mathcal{J}_{1}(h,\beta,z)=\beta^{2}h(x^{\prime\prime})^{2},\qquad\mathcal{J}_{2}(h,\beta,z)=-2\beta h(x^{\prime}),\qquad\mathcal{R}(h,\beta)=\xi\hat{\mathbb{E}}_{1}[\ell_{\alpha}(h,S)].$
We now check the conditions of Lemma G.1.
Boundedness: Note that $|\mathcal{J}_{1}(h,\beta,z)|=|\beta|^{2}|h(x^{\prime\prime})|^{2}\leqslant\beta_{\textnormal{max}}^{2}(BC_{\Phi})^{2}$, where in the second step we used $|\beta|\leqslant\beta_{\textnormal{max}}$ for $\beta\in\mathcal{I}$, and we used $h\in\mathcal{H}_{B}$ and 1 to apply Lemma N.4. Similarly, note that $|\mathcal{J}_{2}(h,\beta,z)|=2|\beta||h(x^{\prime})|\leqslant 2\beta_{\textnormal{max}}BC_{\Phi}$, where in the second step we used $|\beta|\leqslant\beta_{\textnormal{max}}$ for $\beta\in\mathcal{I}$, and we used $h\in\mathcal{H}_{B}$ and 1 to apply Lemma N.4. Thus $|\mathcal{J}_{1}(h,\beta,z)|$ and $|\mathcal{J}_{2}(h,\beta,z)|$ are both bounded by the sum $\beta_{\textnormal{max}}^{2}(BC_{\Phi})^{2}+2\beta_{\textnormal{max}}BC_{\Phi}$.
Convexity: Write $h=\langle\gamma,\Phi\rangle$ for $\gamma\in\mathbb{R}^{d}$. The map $h\mapsto\mathcal{J}_{1}(h,\beta,z)$ can equivalently be written as $\gamma\mapsto\beta^{2}\gamma^{\top}\Phi(x^{\prime\prime})\Phi(x^{\prime\prime})^{\top}\gamma$, a quadratic whose Hessian equals the positive semidefinite matrix $2\beta^{2}\Phi(x^{\prime\prime})\Phi(x^{\prime\prime})^{\top}$. Thus $h\mapsto\mathcal{J}_{1}(h,\beta,z)$ is convex. The map $h\mapsto\mathcal{J}_{2}(h,\beta,z)$ can equivalently be written as $\gamma\mapsto-2\beta\gamma^{\top}\Phi(x^{\prime})$, which is linear, hence convex.
Lipschitzness: Write $h=\langle\gamma,\Phi\rangle$ for $\gamma\in\mathbb{R}^{d}$. The map $h\mapsto\mathcal{J}_{1}(h,\beta,z)$ can equivalently be written as $\gamma\mapsto\beta^{2}\gamma^{\top}\Phi(x^{\prime\prime})\Phi(x^{\prime\prime})^{\top}\gamma$. The gradient of this quadratic is given by $\gamma\mapsto 2\beta^{2}\Phi(x^{\prime\prime})\Phi(x^{\prime\prime})^{\top}\gamma$. The norm of this gradient can be bounded by
$\|2\beta^{2}\Phi(x^{\prime\prime})\Phi(x^{\prime\prime})^{\top}\gamma\|_{2}\leqslant 2|\beta|^{2}\|\Phi(x^{\prime\prime})\|_{2}^{2}\|\gamma\|_{2}\leqslant 2\beta_{\textnormal{max}}^{2}BC_{\Phi}^{2},$
where in the first step we applied the Cauchy-Schwarz inequality, in the second step we used $|\beta|\leqslant\beta_{\textnormal{max}}$ for $\beta\in\mathcal{I}$, $\|\gamma\|_{2}\leqslant B$, and 1. Next, the map $h\mapsto\mathcal{J}_{2}(h,\beta,z)$ can equivalently be written as $\gamma\mapsto-2\beta\gamma^{\top}\Phi(x^{\prime})$. The gradient of this linear map is given by $\gamma\mapsto-2\beta\Phi(x^{\prime})$. The norm of this gradient can be bounded by $2|\beta|\|\Phi(x^{\prime})\|\leqslant 2\beta_{\textnormal{max}}C_{\Phi}$, where we used $|\beta|\leqslant\beta_{\textnormal{max}}$ for $\beta\in\mathcal{I}$ and 1. Thus the norm of each of these gradients is bounded by the sum $\rho_{1}:=2\beta_{\textnormal{max}}^{2}BC_{\Phi}^{2}+2\beta_{\textnormal{max}}C_{\Phi}$, and the maps $h\mapsto\mathcal{J}_{1}(h,\beta,z)$ and $h\mapsto\mathcal{J}_{2}(h,\beta,z)$ are both $\rho_{1}$-Lipschitz.
Strong convexity: Since $h\mapsto\ell_{\alpha}(h,s)$ is convex for all $s\in\mathbb{R}$ by Lemma N.3 and since $h\mapsto\hat{\mathbb{E}}_{2}[\beta h]$ is linear, the map $h\mapsto\xi\hat{\mathbb{E}}_{1}[\ell_{\alpha}(h,S)]-2\hat{\mathbb{E}}_{2}[\beta h]$ is convex. Consider the map $h\mapsto\hat{\mathbb{E}}_{3}[\beta^{2}h^{2}]$. Writing $h=\langle\gamma,\Phi\rangle$ for $\gamma\in\mathbb{R}^{d}$, this can be rewritten as $\gamma\mapsto\beta^{2}\gamma^{\top}\hat{\Sigma}\gamma$, a quadratic whose Hessian equals $2\beta^{2}\hat{\Sigma}$. By $\beta\geqslant\beta_{\textnormal{min}}$ for $\beta\in\mathcal{I}$ and 3, it follows that with probability $1-o(n_{3}^{-1})=1-o(n_{2}^{-1}+n_{3}^{-1})$, the map $h\mapsto\hat{\mathbb{E}}_{2,3}[f_{1}(h,Z)]$ is $\mu_{1}$-strongly convex, where $Z=(X^{\prime\prime},X^{\prime})$ with $X^{\prime}$ is uniform over $\mathcal{X}_{2}$ and $X^{\prime\prime}$ is uniform over $\mathcal{X}_{3}$, and where $\mu_{1}:=2\beta_{\textnormal{min}}^{2}c_{\textnormal{min}}$. In particular, $h\mapsto\frac{1}{\lambda}\hat{L}(h,\beta)$ is convex.
Let $\widetilde{C}_{1}=\frac{4\rho_{1}^{2}}{\mu_{1}}$. Let $K_{1}$ denote the Lipschitz constant of the process $\bar{W}_{\beta}$, where $K_{1}\leqslant K_{1,\lambda}$ with probability tending to unity conditional on $\mathcal{S}_{1}$ by 4 and Lemma H.4. From Lemma G.1 applied with $\xi=1/\lambda$, $\mathcal{L}=\frac{1}{\lambda}\bar{L}$, and ^$\hat{\mathcal{L}}=\frac{1}{\lambda}\hat{L}$, and $W=(\bar{L}-\hat{L})/\lambda$, we obtain that conditional on $\mathcal{S}_{1}$, for sufficiently large $n_{2},n_{3}$, with probability at least $1-\frac{\delta}{4}$, we have for Term (I) from (18),
$\frac{1}{\lambda}\text{Term (I)}\leqslant\sqrt{\frac{16\widetilde{C}_{1}K_{1,\lambda}/\lambda}{\delta/4}\left(\frac{1}{n_{2}}+\frac{1}{n_{3}}\right)}.$
Thus
Term (I) $\leqslant\sqrt{\frac{64\widetilde{C}_{1}\lambda K_{1,\lambda}}{\delta}\left(\frac{1}{n_{2}}+\frac{1}{n_{3}}\right)}=A_{1}\lambda\sqrt{\frac{1}{n_{2}}+\frac{1}{n_{3}}},$
where we define $A_{1}=\sqrt{\frac{64\widetilde{C}_{1}a_{1}}{\delta}}$. Since the right-hand side does not depend on $\mathcal{S}_{1}$, the same bound holds when we uncondition on $\mathcal{S}_{1}$.
Term (II): We work conditional on $\mathcal{S}_{2}$, $\mathcal{S}_{3}$. The risks $\hat{L}$ and $\tilde{L}$ share the same data-independent regularization $\lambda\hat{\mathbb{E}}_{3}[\beta^{2}h^{2}]+\lambda\hat{\mathbb{E}}_{2}[-2\beta h]$. Write $z=(x,s)$ for $x\in\mathcal{X}$ and $s\in[0,1]$. Fixing $\beta$, we apply Lemma G.1 with the objective function $(h,z)\mapsto f(h,z)=\mathcal{J}(h,\beta,z)+\mathcal{R}(h,\beta)$, where
$\mathcal{J}(h,\beta,z)=\ell_{\alpha}(h(x),s),\qquad\mathcal{R}(h,\beta)=\lambda\hat{\mathbb{E}}_{3}[\beta^{2}h^{2}]+\lambda\hat{\mathbb{E}}_{2}[-2\beta h].$
Since the empirical risk $\hat{L}$ is computed over the i.i.d. sample $Z_{i}=(X_{i},S_{i})$ for $i\in[n_{1}]$, we use the modified version of Lemma G.1 given in Remark 1. In particular, we check boundedness, convexity, and Lipschitzness of $\mathcal{J}$ without writing it as a sum $\mathcal{J}_{1}+\mathcal{J}_{2}$.
Boundedness: we have the uniform bound, for all $h,\beta,z$
$|\mathcal{J}(h,\beta,z)|\leqslant(1-\alpha)|h(x)-s|\leqslant(1-\alpha)(|h(x)|+1)\leqslant(1-\alpha)(BC_{\Phi}+1),$ (19)
where in the first step we used Lemma N.1, in the second step we used the triangle inequality and $s\in[0,1]$, and in the third step we used $h\in\mathcal{H}_{B}$ and 1 to apply Lemma N.4.
Convexity: By Lemma N.3, $h\mapsto\mathcal{J}(h,\beta,z)$ is convex.
Lipschitzness: Fix $h=\langle\gamma,\Phi\rangle$ and $h^{\prime}=\langle\gamma^{\prime},\Phi\rangle$ in $\mathcal{H}_{B}$, where $\gamma,\gamma^{\prime}\in\mathbb{R}^{d}$. Note that
$|\mathcal{J}(h,\beta,z)-\mathcal{J}(h,\beta,z)|=|\ell_{\alpha}(h(x),s)-\ell_{\alpha}(h^{\prime}(x),s)|\leqslant(1-\alpha)|h(x)-h^{\prime}(x)|\leqslant(1-\alpha)C_{\Phi}\|h-h^{\prime}\|,$
where in the second step we used Lemma N.2, and in the third step we used 1 to apply Lemma N.4. Thus $h\mapsto\mathcal{J}(h,\beta,z)$ is $\rho_{2}$-Lipschitz, where $\rho_{2}:=(1-\alpha)C_{\Phi}$.
Strong convexity: To analyze $\mathcal{R}$, first observe that since $h\mapsto\lambda\hat{\mathbb{E}}_{2}[-2\beta h]$ is linear, it is convex. Writing $h=\langle\gamma,\Phi\rangle$ for $\gamma\in\mathbb{R}^{d}$, the term $h\mapsto\lambda\hat{\mathbb{E}}_{3}[\beta^{2}h^{2}]$ in $\mathcal{R}$ can be rewritten as $\gamma\mapsto\lambda\beta^{2}\gamma^{\top}\hat{\Sigma}\gamma$, a quadratic whose Hessian equals $2\lambda\beta^{2}\hat{\Sigma}$. By $\beta\geqslant\beta_{\textnormal{min}}$ for $\beta\in\mathcal{I}$ and 3, it follows that with probability $1-o(n_{3}^{-1})$ over $\mathcal{S}_{2},\mathcal{S}_{3}$, the map $h\mapsto\mathcal{R}(h,\beta)$ is $\mu_{2}$-strongly convex, where $\mu_{2}(\lambda):=2\lambda\beta_{\textnormal{min}}^{2}c_{\textnormal{min}}$.
Let $\widetilde{C}_{2}(\lambda)=\frac{4\rho_{2}^{2}}{\mu_{2}(\lambda)}$. Let $K_{2}$ denote the Lipschitz constant of the process $W_{\beta}$; recall that conditional on $\mathcal{S}_{2},\mathcal{S}_{3}$, $K_{2}\leqslant K_{2,\lambda}$ deterministically on the event $C_{1}\leqslant C_{1,\textnormal{max}}$ by Lemma H.4. By the version of Lemma G.1 given in Remark 1, conditional on $\mathcal{S}_{2}$, $\mathcal{S}_{3}$, if $h\mapsto\mathcal{R}(h,\beta)$ is $\mu_{2}(\lambda)$-strongly convex, and if $C_{1}\leqslant C_{1,\textnormal{max}}$, then for sufficiently large $n_{1}$, with probability at least $1-\frac{\delta}{8}$, we have
Term (II) $\leqslant\sqrt{\frac{16\widetilde{C}_{2}(\lambda)K_{2,\lambda}}{(\delta/8)n_{1}}}=\frac{A_{2}}{\sqrt{\lambda n_{1}}},$ (20)
where we define $A_{2}=\sqrt{\frac{128\widehat{C}_{2}a_{2}}{\delta}}$ and $\widehat{C}_{2}=\frac{4\rho_{2}^{2}}{2\beta_{\textnormal{min}}^{2}c_{\textnormal{min}}}$. Unconditioning on $\mathcal{S}_{2},\mathcal{S}_{3}$, since $\mathcal{R}(h,\beta)$ is $\mu_{2}(\lambda)$-strongly convex with probability tending to unity by the above analysis, and since by 3 we have $C_{1}\leqslant C_{1,\textnormal{max}}$ with probability tending to unity, we deduce that for sufficiently large $n_{1},n_{2},n_{3}$, with probability at least $1-\frac{\delta}{4}$, (20) still holds.
Term (III): We work conditional on $\mathcal{S}_{2},\mathcal{S}_{3}$. Since $\tilde{h}$ from (17) lies in $\mathcal{H}_{B}$, we may use the bound in Equation 19 to obtain $\sup_{x\in\mathcal{X}}|\ell_{\alpha}(\tilde{h},S)|\leqslant(1-\alpha)(BC_{\Phi}+1)$. Thus by Hoeffding’s inequality (Hoeffding, , 1963), with probability at least $1-\frac{\delta}{4}$ we have
$(\hat{L}-\tilde{L})(\tilde{h},\tilde{\beta})=(\hat{\mathbb{E}}_{1}-\mathbb{E}_{1})[\ell_{\alpha}(\tilde{h},S)]\leqslant\frac{(1-\alpha)(BC_{\Phi}+1)\sqrt{\frac{1}{2}\log\frac{2}{\delta/4}}}{\sqrt{n_{1}}}.$
[…]
Term (IV): Note that we may write
$(\tilde{L}-L)(h^{*},\beta^{*})$ $=(\hat{\mathbb{E}}_{2}-\mathbb{E}_{2})[\lambda(\beta^{*}h^{*})^{2}]+(\hat{\mathbb{E}}_{3}-\mathbb{E}_{3})[-2\lambda\beta^{*}h^{*}].$
Since $\|h^{*}\|\leqslant B$ by $h^{*}\in\mathcal{H}_{B}$ and since 1 holds, we may apply Lemma N.4 to deduce that $\sup_{x\in\mathcal{X}}|h^{*}(x)|\leqslant BC_{\Phi}$. Consequently, for $\beta\in\mathcal{I}$, we have the uniform bound $\sup_{x\in\mathcal{X}}|\beta h^{*}(x)|\leqslant\beta_{\textnormal{max}}BC_{\Phi}$. By Hoeffding’s inequality (Hoeffding, , 1963), with probability at least $1-\frac{\delta}{8}$, we have
$|(\hat{\mathbb{E}}_{2}-\mathbb{E}_{2})[\lambda(\beta^{*}h^{*})^{2}]|$ $\leqslant\frac{\lambda(\beta_{\textnormal{max}}BC_{\Phi})^{2}\sqrt{\frac{1}{2}\log\frac{2}{\delta/8}}}{\sqrt{n_{2}}}.$
By another application of Hoeffding’s inequality, with probability at least $1-\frac{\delta}{8}$, we have
$|(\hat{\mathbb{E}}_{3}-\mathbb{E}_{3})[-2\lambda\beta^{*}h^{*}]|$ $\leqslant\frac{4\lambda(\beta_{\textnormal{max}}BC_{\Phi})\sqrt{\frac{1}{2}\log\frac{2}{\delta/8}}}{\sqrt{n_{3}}}.$
Summing, with probability at least $1-\delta$ we have the bound
$(\tilde{L}-L)(h^{*},\beta^{*})$ $\leqslant\frac{\lambda(\beta_{\textnormal{max}}BC_{\Phi})^{2}\sqrt{\frac{1}{2}\log\frac{16}{\delta}}}{\sqrt{n_{2}}}+\frac{4\lambda(\beta_{\textnormal{max}}BC_{\Phi})\sqrt{\frac{1}{2}\log\frac{16}{\delta}}}{\sqrt{n_{3}}}.$
[…]
$A_{4}=\sqrt{2}(\beta_{\textnormal{max}}BC_{\Phi})\sqrt{\frac{1}{2}\log\frac{16}{\delta}}\max\left\{\beta_{\textnormal{max}}BC_{\Phi},4\right\}.$
Returning to the analysis of (18), and summing all four terms while defining $A_{5}=A_{1}+A_{4}$, with probability at least $1-\delta$ we obtain a generalization error bound of
$L(\hat{h},\hat{\beta})-L(h^{*},\beta^{*})$ $\leqslant A_{5}\lambda\sqrt{\frac{1}{n_{2}}+\frac{1}{n_{3}}}+A_{3}\frac{1}{\sqrt{n_{1}}}+A_{2}\frac{1}{\sqrt{\lambda}}\frac{1}{\sqrt{n_{1}}}.$
The result follows by taking $c=A_{5}$, $c^{\prime}=A_{3}$, and $c^{\prime\prime}=A_{2}$.
Appendix K Proof of Theorem 4.3
We use the following result to convert the generalization error bound in Theorem 4.2 to a coverage lower bound.
Lemma K.1 (Bounded suboptimality implies bounded gradient for smooth functions).
Let $f:\mathbb{R}^{d^{\prime}}\to\mathbb{R}$, for some positive $d^{\prime}$. Suppose $x^{*}$ is a global minimizer of $f$. Suppose $x^{\prime}$ is such that $f(x^{\prime})\leqslant f(x^{*})+\varepsilon$. Suppose $h\in\mathbb{R}^{d}$ is such that the map $g:\mathbb{R}\to\mathbb{R}$ given by $t\mapsto f(x^{\prime}+th)$ is $L$-smooth, i.e. $|g^{\prime\prime}(h)|$ is uniformly bounded by $L$. Then
$|f^{\prime}(x^{\prime};h)|=|\nabla f(x^{\prime})^{\top}h|\leqslant\sqrt{2L\varepsilon}\|h\|_{2}.$
Proof.
Assume there exists $h$ and $\delta>0$ with $f^{\prime}(x^{\prime};h)>\delta\|h\|$. Setting $y=x^{\prime}-th$,
$f(x^{\prime}-th)\leqslant f(x^{\prime})-tf^{\prime}(x^{\prime};h)+\tfrac{L}{2}t^{2}\|h\|^{2}.$
Set $t=\delta/(L\|h\|)$ to obtain
$f(x^{\prime}-th)\leqslant f(x^{\prime})-\tfrac{\delta^{2}}{L}+\tfrac{\delta^{2}}{2L}=f(x^{\prime})-\tfrac{\delta^{2}}{2L}.$
Since $f(x^{\prime})\leqslant f(x^{*})+\varepsilon$, we have $f(x^{\prime}-th)\leqslant f(x^{*})+\varepsilon-\tfrac{\delta^{2}}{2L}$. If $\delta>\sqrt{2L\varepsilon}$, then $f(x^{\prime}-th)<f(x^{*})$, a contradiction.
A similar argument with $f^{\prime}(x^{\prime};h)<-\delta\|h\|$ and $y=x^{\prime}+th$ yields the same contradiction. Hence $-\sqrt{2L\varepsilon}\|h\|\leqslant f^{\prime}(x^{\prime};h)\leqslant\sqrt{2L\varepsilon}\|h\|$. ∎
By 1 and 5, we may apply Lemma N.5 to deduce that the Hessian of our population risk $L$ from (2) in the basis $\{\phi_{1},\ldots,\phi_{d}\}$ is the block matrix
$\nabla^{2}L(h,\beta)$ $=\begin{bmatrix}\mathbb{E}_{1}[\Phi\Phi^{\top}(f_{S|X}(h)+2\lambda\beta^{2})]&\mathbb{E}_{1}[2\lambda\Phi^{\top}(2\beta h-r)]\\ \mathbb{E}_{1}[2\lambda\Phi(2\beta h-r)]&\mathbb{E}_{1}[2\lambda h^{2}]\end{bmatrix}.$
Thus by $\beta\leqslant\beta_{\textnormal{max}}$, $\|h\|\leqslant B$ for $h\in\mathcal{H}_{B}$, 5, and Jensen’s inequality, we have the uniform bounds
$\sup_{h\in\mathcal{H}_{B},\beta\in\mathbb{R}}|\partial_{\beta}^{2}L(h,\beta)|$ $\leqslant 2\lambda\mathbb{E}_{1}[h^{2}]\leqslant 2\lambda B^{2}\lambda_{\textnormal{max}}(\Sigma)=:\nu_{1}$
and
$\sup_{h\in\mathcal{H},\beta\in\mathcal{I}}\|\nabla_{h}^{2}L(h,\beta)\|_{2}$ $=\|\mathbb{E}_{1}[\Phi\Phi^{\top}(f_{S|X}(h)+2\lambda\beta^{2})]\|_{2}\ \leqslant(C_{f}+2\lambda\beta_{\textnormal{max}}^{2})\lambda_{\textnormal{max}}(\Sigma)=:\nu_{2}.$
By Lemma L.3 and Lemma L.4, a global minimizer of the objective in Equation LR-QR exists, and since $\beta_{\textnormal{min}}\leqslant\beta_{\textnormal{lower}}$, $\beta_{\textnormal{max}}\geqslant\beta_{\textnormal{upper}}$, and $B\geqslant B_{\textnormal{upper}}$, any such minimizer lies in the interior of $\mathcal{H}_{B}\times\mathcal{I}$. Thus we may apply Lemma K.1 to the objective function $L$. We utilize two directional derivatives in the space $\mathcal{H}\times\mathbb{R}$. The first is in the direction $0_{\mathcal{H}}\times 1$, the unit vector in the $\beta$ coordinate. Since $(\hat{h},\hat{\beta})\in\mathcal{H}_{B}\times\mathcal{I}$, the magnitude of the second derivative of $L$ along this direction is bounded by $\nu_{1}$.
The second is in the direction of the vector $r_{\mathcal{H}_{B}}\times 0$, where $r_{\mathcal{H}_{B}}$ the projection of $r$ onto the closed convex set $\mathcal{H}_{B}$ in the Hilbert space induced by the inner product $\langle f,g\rangle=\mathbb{E}_{1}[fg]$. Since $(\hat{h},\hat{\beta})\in\mathcal{H}_{B}\times\mathcal{I}$, the magnitude of the second derivative of $L$ along this direction is bounded by $\nu_{2}$.
Given $\hat{h}$, let $\widehat{\textnormal{Cover}}(X):=\mathbb{P}\left[S\leqslant\hat{h}(X)|X\right]-(1-\alpha)$. Now, on the event $E$ that $L(\hat{h},\hat{\beta})-L(h^{*},\beta^{*})\leqslant\mathcal{E}_{\textnormal{gen}}$, we apply Lemma K.1 with $f$ being $(\gamma,\beta)\mapsto L(h_{\gamma},\beta)$, $x^{*}$ being $(h^{*},\beta^{*})$, $x^{\prime}$ being $(\hat{h},\hat{\beta})$, $\varepsilon=\mathcal{E}_{\textnormal{gen}}$, and the directions specified above, with their respective smoothness parameters derived above. Using the formulas for $\nabla L$ from Lemma N.5 and the bound $\|r_{\mathcal{H}_{B}}\|\leqslant B$, we obtain that on the event $E$,
$|2\lambda\mathbb{E}_{1}[\hat{h}(\hat{\beta}\hat{h}-r)]|\leqslant\mathcal{E}_{1},\qquad|\mathbb{E}_{1}[r_{\mathcal{H}_{B}}\widehat{\textnormal{Cover}}]+\lambda\mathbb{E}_{1}[2\beta r_{\mathcal{H}_{B}}(\hat{\beta}\hat{h}-r)]|\leqslant\mathcal{E}_{2},$
where $\mathcal{E}_{1}=\sqrt{2\nu_{1}\mathcal{E}_{\textnormal{gen}}}$, $\mathcal{E}_{2}=\sqrt{2B^{2}\nu_{2}\mathcal{E}_{\textnormal{gen}}}$.
For any $h$ and $\beta$, we may write
$\mathbb{E}_{1}[r_{\mathcal{H}_{B}}\widehat{\textnormal{Cover}}]$ $=(\mathbb{E}_{1}[r_{\mathcal{H}_{B}}\widehat{\textnormal{Cover}}]+\lambda\mathbb{E}_{1}[2\beta r_{\mathcal{H}_{B}}(\beta h-r)])-\lambda\mathbb{E}_{1}[2\beta(\beta h)(\beta h-r)]-\lambda\mathbb{E}_{1}[2\beta(r_{\mathcal{H}_{B}}-\beta h)(\beta h-r)].$
Evaluating at $(\hat{h},\hat{\beta})$, the first term is at most $\mathcal{E}_{2}$ in magnitude, the second term is at most $\hat{\beta}^{2}\mathcal{E}_{1}$ in magnitude, and the third term equals $2\hat{\beta}\lambda\mathbb{E}_{1}[(r_{\mathcal{H}_{B}}-\hat{\beta}\hat{h})^{2}]$. We deduce
$\mathbb{E}_{1}[r_{\mathcal{H}_{B}}\widehat{\textnormal{Cover}}]\geqslant 2\hat{\beta}\lambda\mathbb{E}_{1}[(r_{\mathcal{H}_{B}}-\hat{\beta}\hat{h})^{2}]-\hat{\beta}^{2}\mathcal{E}_{1}-\mathcal{E}_{2}.$
Since $\widehat{\textnormal{Cover}}\in[-(1-\alpha),\alpha]$,
$|\mathbb{E}_{1}[r\widehat{\textnormal{Cover}}]-\mathbb{E}_{1}[r_{\mathcal{H}_{B}}\widehat{\textnormal{Cover}}]|\leqslant(1-\alpha)\mathbb{E}_{1}[|r-r_{\mathcal{H}_{B}}|].$
We deduce that
$\mathbb{E}_{1}[r\widehat{\textnormal{Cover}}]$ $\geqslant 2\hat{\beta}\lambda\mathbb{E}_{1}[(r_{\mathcal{H}_{B}}-\hat{\beta}\hat{h})^{2}]-\hat{\beta}^{2}\mathcal{E}_{1}-\mathcal{E}_{2}-(1-\alpha)\mathbb{E}_{1}[|r-r_{\mathcal{H}_{B}}|].$
We now bound the quantity $\hat{\beta}^{2}\mathcal{E}_{1}+\mathcal{E}_{2}$. First, since $\sqrt{a+b}\leqslant\sqrt{a}+\sqrt{b}$ for all $a,b\geqslant 0$, Theorem 4.2 implies that
$\sqrt{\mathcal{E}_{\textnormal{gen}}}$ $\leqslant A_{5}^{1/2}\lambda^{1/2}\left(\frac{1}{n_{2}}+\frac{1}{n_{3}}\right)^{1/4}+\frac{A_{3}^{1/2}}{n_{1}^{1/4}}+A_{2}^{1/2}\frac{1}{\lambda^{1/4}}\frac{1}{n_{1}^{1/4}}.$
We may write $\mathcal{E}_{1}=\sqrt{2\nu_{1}\mathcal{E}_{\textnormal{gen}}}=\sqrt{4B^{2}\lambda_{\textnormal{max}}(\Sigma)}\cdot\lambda^{1/2}\sqrt{\mathcal{E}_{\textnormal{gen}}}$, so that for $\hat{\beta}\in\mathcal{I}$ we have
$\hat{\beta}^{2}\mathcal{E}_{1}\leqslant\beta_{\textnormal{max}}^{2}\sqrt{4B^{2}\lambda_{\textnormal{max}}(\Sigma)}\cdot\lambda^{1/2}\sqrt{\mathcal{E}_{\textnormal{gen}}}=:A_{6}\lambda^{1/2}\sqrt{\mathcal{E}_{\textnormal{gen}}}.$
Using the inequality $\sqrt{a+b}\leqslant\sqrt{a}+\sqrt{b}$ for all $a,b\geqslant 0$, we may bound
$\mathcal{E}_{2}$ $=\sqrt{2B^{2}\nu_{2}\mathcal{E}_{\textnormal{gen}}}\leqslant\sqrt{4B^{2}\beta_{\textnormal{max}}^{2}\lambda_{\textnormal{max}}(\Sigma)}\cdot\lambda^{1/2}\sqrt{\mathcal{E}_{\textnormal{gen}}}+\sqrt{2B^{2}C_{f}\lambda_{\textnormal{max}}(\Sigma)}\cdot\sqrt{\mathcal{E}_{\textnormal{gen}}}$
$=:A_{7}\lambda^{1/2}\sqrt{\mathcal{E}_{\textnormal{gen}}}+A_{8}\sqrt{\mathcal{E}_{\textnormal{gen}}},$
Thus
$\hat{\beta}^{2}\mathcal{E}_{1}+\mathcal{E}_{2}$ $\leqslant A_{6}\lambda^{1/2}\sqrt{\mathcal{E}_{\textnormal{gen}}}+A_{7}\lambda^{1/2}\sqrt{\mathcal{E}_{\textnormal{gen}}}+A_{8}\sqrt{\mathcal{E}_{\textnormal{gen}}}$
$=:A_{9}\lambda^{1/2}\sqrt{\mathcal{E}_{\textnormal{gen}}}+A_{8}\sqrt{\mathcal{E}_{\textnormal{gen}}}.$
Plugging in our bound on $\sqrt{\mathcal{E}_{\textnormal{gen}}}$ and grouping terms according to the power of $\lambda$, we deduce that $\hat{\beta}^{2}\mathcal{E}_{1}+\mathcal{E}_{2}\leqslant\mathcal{E}_{\textnormal{cov}}$, where $\mathcal{E}_{\textnormal{cov}}$ equals
$A_{10}\left(\frac{1}{n_{2}}+\frac{1}{n_{3}}\right)^{1/4}\lambda+A_{11}\left(\frac{1}{n_{1}^{1/4}}+\left(\frac{1}{n_{2}}+\frac{1}{n_{3}}\right)^{1/4}\right)\lambda^{1/2}+A_{12}\frac{1}{n_{1}^{1/4}}\lambda^{1/4}+A_{13}\frac{1}{n_{1}^{1/4}}+A_{14}\frac{1}{n_{1}^{1/4}}\lambda^{-1/4}$
and where $A_{10},\ldots,A_{14}$ are the positive constants given in Appendix F. It follows that on the event $E$,
$\mathbb{E}_{1}[r\widehat{\textnormal{Cover}}]\geqslant(1-\alpha)+2\hat{\beta}\lambda\mathbb{E}_{1}[(r_{\mathcal{H}_{B}}-\hat{\beta}\hat{h})^{2}]-\mathcal{E}_{\textnormal{cov}}-(1-\alpha)\mathbb{E}_{1}[|r-r_{\mathcal{H}_{B}}|].$
By Theorem 4.2, $E$ occurs with probability $1-\delta$ for sufficiently large $n_{1},n_{2},n_{3}$, and we may conclude.
Appendix L Unconstrained existence and boundedness
In this section, we prove apriori existence and boundedness of unconstrained global minimizers of the population objective Equation LR-QR. We write $(h_{\lambda}^{*},\beta_{\lambda}^{*})$ for a minimizer of the unconstrained objective in Equation LR-QR with regularization strength $\lambda\geqslant 0$.
In Lemma L.1, we show that under 10, we may eliminate $\beta$ from Equation LR-QR, so that Equation LR-QR is equivalent to solving the following unconstrained optimization problem over $h$:
$\min_{h\in\mathcal{H}\setminus\{0\}}\mathbb{E}_{1}[\ell_{\alpha}(h,S)]-\lambda\frac{\mathbb{E}_{1}[rh]^{2}}{\mathbb{E}_{1}[h^{2}]}.$ (21)
Lemma L.1.
Under 10, for $\lambda\geqslant 0$, given any minimizer $(h_{\lambda}^{*},\beta_{\lambda}^{*})$ of the objective in Equation LR-QR with regularization $\lambda$, $h_{\lambda}^{*}$ is a minimizer of the objective in Equation 21 with regularization $\lambda$. Conversely, if $h$ is a minimizer of the objective in Equation 21 with regularization $\lambda$, then there exists a minimizer $(h_{\lambda}^{*},\beta_{\lambda}^{*})$ of the objective in Equation LR-QR with regularization $\lambda$ such that $h_{\lambda}^{*}=h$.
Proof.
By 10, the minimization in Equation LR-QR with regularization $\lambda$ can be taken over $\mathcal{H}\setminus\{0\}$. Further, since the projection of $r$ onto $\textnormal{span}\{h\}:=\{ch:c\in\mathbb{R}\}$, for $h\neq 0$ is given by $\frac{\mathbb{E}_{1}[rh]}{\mathbb{E}_{1}[h^{2}]}h$, we may explicitly minimize the objective in Equation LR-QR over $\beta$ via
$\ell_{\alpha}(h,S)+\lambda\min_{\beta\in\mathbb{R}}\mathbb{E}_{1}[(\beta h-r)^{2}]=\ell_{\alpha}(h,S)+\lambda\mathbb{E}_{1}\left[\left(\frac{\mathbb{E}_{1}[rh]}{\mathbb{E}_{1}[h^{2}]}h-r\right)^{2}\right]$
$=\ell_{\alpha}(h,S)+\lambda\left(\mathbb{E}_{1}[r^{2}]-\mathbb{E}_{1}\left[\left(\frac{\mathbb{E}_{1}[rh]}{\mathbb{E}_{1}[h^{2}]}h\right)^{2}\right]\right)=\ell_{\alpha}(h,S)+\lambda\left(\mathbb{E}_{1}[r^{2}]-\frac{\mathbb{E}_{1}[rh]^{2}}{\mathbb{E}_{1}[h^{2}]}\right),$
where in the second step we applied the Pythagorean theorem. Since the term $\lambda\mathbb{E}_{1}[r^{2}]$ does not depend on the optimization variable $h$, we may drop it from the objective, which yields the objective in Equation 21. It follows that $h$ is a minimizer of the objective in Equation 21 iff $h=h_{\lambda}^{*}$ for some minimizer $(h_{\lambda}^{*},\beta_{\lambda}^{*})$ of the objective of Equation LR-QR. ∎
Lemma L.2.
Let $r_{\mathcal{H}}$ denote the projection of $r$ onto $\mathcal{H}$ in the Hilbert space induced by the inner product $\langle f,g\rangle=\mathbb{E}_{1}[fg]$. Then under 5 and 9, there exists $\theta^{*}>0$ such that $\mathbb{E}_{1}[S]-\alpha^{-1}\mathbb{E}_{1}[\ell_{\alpha}(\theta^{*}r_{\mathcal{H}},S)]>0$.
Proof.
Define $g:\mathbb{R}\to\mathbb{R}$ by $g(\theta)=\mathbb{E}_{1}[S]-\alpha^{-1}\mathbb{E}_{1}[\ell_{\alpha}(\theta^{*}r_{\mathcal{H}},S)]$. Clearly $g(0)=0$. Note that by 5, $\mathbb{P}_{S|X}[S=0]=0$, so that
$g^{\prime}(0)=-\alpha^{-1}\mathbb{E}_{1}[r_{\mathcal{H}}(\mathbb{P}_{S|X}[S\leqslant 0]-(1-\alpha))]=\alpha^{-1}(1-\alpha)\mathbb{E}_{1}[r_{\mathcal{H}}].$
By 9, $\mathbb{E}_{1}[r_{\mathcal{H}}]=\mathbb{E}_{1}[r_{\mathcal{H}}\cdot 1]=\mathbb{E}_{1}[r\cdot 1]=\mathbb{E}_{1}[r]=1$, so $g^{\prime}(0)>0$. Thus there exists $\theta^{*}>0$ such that $g(\theta^{*})>g(0)=0$, as claimed. ∎
Lemma L.3 (Existence of unconstrained minimizers).
Under 2, 5, 6, 7, 8, 9, and 10, for each $\lambda\geqslant 0$, there exists a global minimizer $(h_{\lambda}^{*},\beta_{\lambda}^{*})$ of the objective in Equation LR-QR.
Proof.
Fix $\lambda\geqslant 0$. By 10 and Lemma L.1, it suffices to show that there exists a global minimizer of the objective in Equation 21. Let $G(h)$ denote the objective of Equation 21. Define the function $\tilde{h}=\theta^{*}r_{\mathcal{H}}\in\mathcal{H}\setminus\{0\}$, where $\theta^{*}$ is chosen to satisfy Lemma L.2. With $c_{\textnormal{indep}}$ from 6, define $\tilde{B}(\lambda):=2c_{\textnormal{indep}}^{-1}(1+\alpha^{-1}\mathbb{E}_{1}[\ell_{\alpha}(\tilde{h},S)])>0$ and
$\tilde{b}(\lambda):=\frac{1}{2}\lambda_{\textnormal{max}}(\Sigma)^{-1/2}(\mathbb{E}_{1}[S]-\alpha^{-1}\mathbb{E}_{1}[\ell_{\alpha}(\tilde{h},S)])>0.$
We show that if $\|h\|\geqslant\tilde{B}(\lambda)$ or $\|h\|\leqslant\tilde{b}(\lambda)$, then $G(h)>G(\tilde{h})$. Consequently, the minimization in Equation 21 can be taken over the compact set $\{\langle\gamma,\Phi\rangle:\tilde{b}(\lambda)\leqslant\|\gamma\|_{2}\leqslant\tilde{B}(\lambda)\}\subseteq\mathcal{H}$, so that by continuity of $G$ on $\mathcal{H}\setminus\{0\}$, a global minimizer $h^{*}_{\lambda}$ exists.
To see this, first suppose $\|h\|\geqslant\tilde{B}(\lambda)$. Then writing $h=\langle\gamma,\Phi\rangle$ for $\gamma\in\mathbb{R}^{d}$ and applying Lemma N.1, the triangle inequality, and $S\in[0,1]$,
$\mathbb{E}_{1}[\ell_{\alpha}(h,S)]\geqslant\alpha\mathbb{E}_{1}[|h-S|]\geqslant\alpha(\mathbb{E}_{1}[|h|]-\mathbb{E}_{1}[|S|])\geqslant\alpha(\mathbb{E}_{1}[|\langle\gamma,\Phi\rangle|]-1).$ (22)
By 6 and our assumption that $\|h\|\geqslant\tilde{B}(\lambda)$, this implies that $\mathbb{E}_{1}[\ell_{\alpha}(h,S)]\geqslant\alpha(\tilde{B}(\lambda)c_{\textnormal{indep}}-1)$. Further, by the Cauchy-Schwarz inequality,
$\frac{\mathbb{E}_{1}[rh]^{2}}{\mathbb{E}_{1}[h^{2}]}\leqslant\sup_{\tilde{h}^{\prime}\in\mathcal{H}\setminus\{0\}}\frac{\mathbb{E}_{1}[r\tilde{h}^{\prime}]^{2}}{\mathbb{E}_{1}[(\tilde{h}^{\prime})^{2}]}\leqslant\mathbb{E}_{1}[r_{\mathcal{H}}^{2}].$
Thus by Lemma N.1 and 8, $G(h)\geqslant\alpha(\tilde{B}(\lambda)c_{\textnormal{indep}}-1)-\lambda\mathbb{E}_{1}[r_{\mathcal{H}}^{2}]$. To prove the inequality $G(h)>G(\tilde{h})$, it suffices to show that
$\alpha(\tilde{B}(\lambda)c_{\textnormal{indep}}-1)-\lambda\mathbb{E}_{1}[r_{\mathcal{H}}^{2}]>\mathbb{E}_{1}[\ell_{\alpha}(\tilde{h},S)]-\lambda\frac{\mathbb{E}_{1}[r\tilde{h}]^{2}}{\mathbb{E}_{1}[\tilde{h}^{2}]}.$
Indeed, since $\tilde{h}$ is a scalar multiple of $r_{\mathcal{H}}$, we have $\mathbb{E}_{1}[r_{\mathcal{H}}^{2}]=\frac{\mathbb{E}_{1}[r\tilde{h}]^{2}}{\mathbb{E}_{1}[\tilde{h}^{2}]}$, so the inequality reduces to $\alpha(\tilde{B}(\lambda)c_{\textnormal{indep}}-1)>\mathbb{E}_{1}[\ell_{\alpha}(\tilde{h},S)]$. This holds by our choice of $\tilde{B}(\lambda)$, which finishes the argument in this case.
Next, suppose $\|h\|\leqslant\tilde{b}(\lambda)$. By Lemma N.1, the triangle inequality, and $S\in[0,1]$,
$\mathbb{E}_{1}[\ell_{\alpha}(h,S)]\geqslant\alpha\mathbb{E}_{1}[|h-S|]\geqslant\alpha(\mathbb{E}_{1}[S]-\mathbb{E}_{1}[|h|]).$ (23)
As above, the Cauchy-Schwarz inequality implies the bound $\frac{\mathbb{E}_{1}[rh]^{2}}{\mathbb{E}_{1}[h^{2}]}\leqslant\mathbb{E}_{1}[r_{\mathcal{H}}^{2}]$. We deduce that
$G(h)$ $\geqslant\alpha(\mathbb{E}_{1}[S]-\mathbb{E}_{1}[|h|])-\lambda\mathbb{E}_{1}[r_{\mathcal{H}}^{2}].$
Writing $h=\langle\gamma,\Phi\rangle$ for $\gamma\in\mathbb{R}^{d}$, our assumption that $\|h\|\leqslant\tilde{b}(\lambda)$ implies that
$\mathbb{E}_{1}[|h|]\leqslant\mathbb{E}_{1}[|h|^{2}]^{1/2}=\mathbb{E}_{1}[\gamma^{\top}\Phi\Phi^{\top}\gamma]^{1/2}\leqslant\tilde{b}(\lambda)\lambda_{\textnormal{max}}(\Sigma)^{1/2},$
which when plugged into our lower bound on $G(h)$ yields
$G(h)\geqslant\alpha(\mathbb{E}_{1}[S]-\tilde{b}(\lambda)\lambda_{\textnormal{max}}(\Sigma)^{1/2})-\lambda\mathbb{E}_{1}[r_{\mathcal{H}}^{2}].$
To prove the inequality $G(h)>G(\tilde{h})$, it suffices to show that
$\alpha(\mathbb{E}_{1}[S]-\tilde{b}(\lambda)\lambda_{\textnormal{max}}(\Sigma)^{1/2})-\lambda\mathbb{E}_{1}[r_{\mathcal{H}}^{2}]>\mathbb{E}_{1}[\ell_{\alpha}(\tilde{h},S)]-\lambda\frac{\mathbb{E}_{1}[r\tilde{h}]^{2}}{\mathbb{E}_{1}[\tilde{h}^{2}]}.$
As above, since $\tilde{h}$ is a scalar multiple of $r_{\mathcal{H}}$, we have $\mathbb{E}_{1}[r_{\mathcal{H}}^{2}]=\frac{\mathbb{E}_{1}[r\tilde{h}]^{2}}{\mathbb{E}_{1}[\tilde{h}^{2}]}$, so the inequality reduces to
$\alpha(\mathbb{E}_{1}[S]-\tilde{b}(\lambda)\lambda_{\textnormal{max}}(\Sigma)^{1/2})>\mathbb{E}_{1}[\ell_{\alpha}(\tilde{h},S)].$
This holds for our choice of $\tilde{b}(\lambda)$, finishing the proof. ∎
Lemma L.4 (Bounds on unconstrained minimizers).
Under the conditions used in Lemma L.3, for all $\lambda>0$, for any minimizer $(h_{\lambda}^{*},\beta_{\lambda}^{*})$ of the objective in Equation LR-QR, we have that $\|h_{\lambda}^{*}\|\in(B_{\textnormal{lower}},B_{\textnormal{upper}})$ and $\beta_{\lambda}^{*}\in(\beta_{\textnormal{lower}},\beta_{\textnormal{upper}})$, where
$B_{\textnormal{lower}}$ $=\frac{1}{2}\lambda_{\textnormal{max}}(\Sigma)^{-1/2}(\mathbb{E}_{1}[S]-\alpha^{-1}\mathbb{E}_{1}[\ell_{\alpha}(\theta^{*}r_{\mathcal{H}},S)])>0,$ (24)
$B_{\textnormal{upper}}$ $=2c_{\textnormal{indep}}^{-1}(\alpha^{-1}\mathbb{E}_{1}[\ell_{\alpha}(\theta^{*}r_{\mathcal{H}},S)]+1),\qquad\beta_{\textnormal{lower}}=\frac{c_{\textnormal{align}}}{B_{\textnormal{upper}}\lambda_{\textnormal{max}}(\Sigma)^{1/2}}>0,\qquad\beta_{\textnormal{upper}}=\frac{\mathbb{E}_{1}[r^{2}]^{1/2}}{B_{\textnormal{lower}}\lambda_{\textnormal{min}}(\Sigma)^{1/2}},$
and where $\theta^{*}>0$ is as in Lemma L.2 and $r_{\mathcal{H}}$ denotes the projection of $r$ onto $\mathcal{H}$ in the Hilbert space induced by the inner product $\langle f,g\rangle=\mathbb{E}_{1}[fg]$.
Proof.
In order to derive our bounds, we consider the reparametrized optimization problem
$\min_{h\in\mathcal{H}\setminus\{0\}}\xi\mathbb{E}_{1}[\ell_{\alpha}(h,S)]-\frac{\mathbb{E}_{1}[rh]^{2}}{\mathbb{E}_{1}[h^{2}]}$ (25)
for $\xi\geqslant 0$. We claim that for $\xi>0$, any minimizer of the objective in Equation 25 is of the form $h_{1/\xi}^{*}$. To see this, note that for $\xi>0$, the objective of Equation 21 with regularization $\lambda=1/\xi$ can be obtained by scaling the objective of Equation 25 by the positive factor $1/\xi$. Next, by 10, we may apply Lemma L.1 to deduce that $h\in\mathcal{H}\setminus\{0\}$ is a minimizer of the objective in Equation 21 with regularization $\lambda=1/\xi$ iff $h=h_{1/\xi}^{*}$.
In particular, by Lemma L.3, for all $\xi>0$, there exists a global minimizer of Equation 25 with regularization $\xi$. In the case that $\xi=0$, it is clear that any minimizer $h^{*}_{\infty}$ of the objective in Equation 25 with regularization $\xi=0$ has the form $h^{*}_{\infty}=\theta r_{\mathcal{H}}$ for some scalar $\theta>0$.
Since there exists a minimizer of the objective in Equation 25 for all regularizations $\xi$ in the interval $[0,\infty)$, we may apply Lemma M.1 to deduce that for all $\xi>0$ we have $\mathbb{E}_{1}[\ell_{\alpha}(h_{1/\xi}^{*},S)]\leqslant\mathbb{E}_{1}[\ell_{\alpha}(h_{\infty}^{*},S)]$.
We prove lower and upper bounds on $\|h_{1/\xi}^{*}\|$ for all $\xi>0$. We begin with the lower bound.
Lower bound: By (23), we have $\mathbb{E}_{1}[\ell_{\alpha}(h_{1/\xi}^{*},S)]\geqslant\alpha(\mathbb{E}_{1}[S]-\mathbb{E}_{1}[|h_{1/\xi}^{*}|]).$ Rearranging, we obtain the lower bound
$\mathbb{E}_{1}[|h_{1/\xi}^{*}|]\geqslant\mathbb{E}_{1}[S]-\alpha^{-1}\mathbb{E}_{1}[\ell_{\alpha}(h_{\infty}^{*},S)].$
By Lemma L.2, there exists $\theta^{*}>0$ such that $\mathbb{E}_{1}[S]-\alpha^{-1}\mathbb{E}_{1}[\ell_{\alpha}(\theta^{*}r_{\mathcal{H}},S)]>0$. Setting $h_{\infty}^{*}=\theta^{*}r_{\mathcal{H}}$ and plugging in the expression for $B_{\textnormal{lower}}$ given in (24), our lower bound becomes $\mathbb{E}_{1}[|h_{1/\xi}^{*}|]>\lambda_{\textnormal{max}}(\Sigma)^{1/2}B_{\textnormal{lower}}$. We now convert this $L^{1}$ norm bound to an $L^{2}$ norm bound as follows. Write $h_{1/\xi}^{*}=\langle\gamma_{1/\xi}^{*},\Phi\rangle$ for $\gamma_{1/\xi}^{*}\in\mathbb{R}^{d}$. By the Cauchy-Schwarz inequality, we obtain the upper bound
$\mathbb{E}_{1}[|h_{1/\xi}^{*}|]\leqslant\mathbb{E}_{1}[|h_{1/\xi}^{*}|^{2}]^{1/2}=\mathbb{E}_{1}[(\gamma_{1/\xi}^{*})^{\top}\Phi\Phi^{\top}\gamma_{1/\xi}^{*}]^{1/2}\leqslant\lambda_{\textnormal{max}}(\Sigma)^{1/2}\|\gamma_{1/\xi}^{*}\|_{2}.$
Combining this with the lower bound $\mathbb{E}_{1}[|h_{1/\xi}^{*}|]>\lambda_{\textnormal{max}}(\Sigma)^{1/2}B_{\textnormal{lower}}$, we deduce that $\|h_{1/\xi}^{*}\|=\|\gamma_{1/\xi}^{*}\|_{2}>B_{\textnormal{lower}}$, as claimed.
Upper bound: We prove the upper bound in a similar manner. By the first two steps in (22), and using $S\in[0,1]$, we have
$\mathbb{E}_{1}[\ell_{\alpha}(h_{1/\xi}^{*},S)]\geqslant\alpha(\mathbb{E}_{1}[|h_{1/\xi}^{*}|]-\mathbb{E}_{1}[|S|])\geqslant\alpha(\mathbb{E}_{1}[|h_{1/\xi}^{*}|]-1).$
Rearranging, we obtain the upper bound $\mathbb{E}_{1}[|h_{1/\xi}^{*}|]\leqslant\alpha^{-1}\mathbb{E}_{1}[\ell_{\alpha}(h_{\infty}^{*},S)]+1$. Write $h_{1/\xi}^{*}=\langle\gamma_{1/\xi}^{*},\Phi\rangle$ for $\gamma_{1/\xi}^{*}\in\mathbb{R}^{d}$. Since we have already established that $\|h_{1/\xi}^{*}\|>B_{\textnormal{lower}}>0$, we know that $\gamma_{1/\xi}^{*}\neq 0$. Thus we may write
$\mathbb{E}_{1}[|h_{1/\xi}^{*}|]$ $=\mathbb{E}_{1}[|\langle\gamma_{1/\xi}^{*},\Phi\rangle|]=\|\gamma_{1/\xi}^{*}\|_{2}\mathbb{E}_{1}\left[\left|\left\langle\frac{\gamma_{1/\xi}^{*}}{\|\gamma_{1/\xi}^{*}\|_{2}},\Phi\right\rangle\right|\right].$
By 6, this is at least $\|\gamma_{1/\xi}^{*}\|_{2}c_{\textnormal{indep}}$. Combining these upper and lower bounds on $\mathbb{E}_{1}[|h_{1/\xi}^{*}|]$, we obtain $\|\gamma_{1/\xi}^{*}\|_{2}c_{\textnormal{indep}}\leqslant\alpha^{-1}\mathbb{E}_{1}[\ell_{\alpha}(h_{\infty}^{*},S)]+1$. Isolating $\|\gamma_{1/\xi}^{*}\|_{2}$, we have
$\|h^{*}_{1/\xi}\|=\|\gamma_{1/\xi}^{*}\|_{2}\leqslant c_{\textnormal{indep}}^{-1}(\alpha^{-1}\mathbb{E}_{1}[\ell_{\alpha}(h_{\infty}^{*},S)]+1)<B_{\textnormal{upper}},$
as claimed.
Having established $0<B_{\textnormal{lower}}<\inf_{\lambda>0}\|h_{\lambda}^{*}\|\leqslant\sup_{\lambda>0}\|h_{\lambda}^{*}\|<B_{\textnormal{upper}}<\infty$, we turn to upper and lower bounds on $\beta^{*}_{\lambda}$. As shown in the proof of Lemma L.1, if $(h_{\lambda}^{*},\beta_{\lambda}^{*})$ is a minimizer of the objective in Equation LR-QR with regularization $\lambda$, then $\beta_{\lambda}^{*}=\frac{\mathbb{E}_{1}[rh_{\lambda}^{*}]}{\mathbb{E}_{1}[|h_{\lambda}^{*}|^{2}]}$. By 7, $\frac{\mathbb{E}_{1}[rh_{0}^{*}]}{\mathbb{E}_{1}[|h_{0}^{*}|^{2}]^{1/2}}\geqslant c_{\textnormal{align}}>0$ for some minimizer $(h_{0}^{*},\beta_{0}^{*})$ of the objective in Equation LR-QR with regularization 00. By 10 and Lemma L.1, $h$ is a minimizer of the objective in Equation 21 with regularization $\lambda\geqslant 0$ iff $h=h_{\lambda}^{*}$ for some minimizer $(h_{\lambda}^{*},\beta_{\lambda}^{*})$ of the objective in Equation LR-QR. Thus by Lemma L.3, for all $\lambda\geqslant 0$, there exists a global minimizer of Equation 21, and we may apply Lemma M.1 to Equation 21 to deduce that for any $\lambda\geqslant 0$ we have $\frac{\mathbb{E}_{1}[rh_{\lambda}^{*}]}{\mathbb{E}_{1}[|h_{\lambda}^{*}|^{2}]^{1/2}}\geqslant c_{\textnormal{align}}>0$. Consequently, by our bounds on $h_{\lambda}^{*}$, 8, and the Cauchy-Schwarz inequality, if we write $h_{\lambda}^{*}=\langle\gamma_{\lambda}^{*},\Phi\rangle$ for $\gamma_{\lambda}^{*}\in\mathbb{R}^{d}$, then we have
$\beta_{\lambda}^{*}\geqslant\frac{c_{\textnormal{align}}}{\mathbb{E}_{1}[|h_{\lambda}^{*}|^{2}]^{1/2}}=\frac{c_{\textnormal{align}}}{\mathbb{E}_{1}[(\gamma_{\lambda}^{*})^{\top}\Phi\Phi^{\top}\gamma_{\lambda}^{*}]^{1/2}}>\frac{c_{\textnormal{align}}}{B_{\textnormal{upper}}\lambda_{\textnormal{max}}(\Sigma)^{1/2}}=:\beta_{\textnormal{lower}}$
and
$\beta_{\lambda}^{*}\leqslant\frac{\mathbb{E}_{1}[r^{2}]^{1/2}}{\mathbb{E}_{1}[|h_{\lambda}^{*}|^{2}]^{1/2}}<\frac{\mathbb{E}_{1}[r^{2}]^{1/2}}{B_{\textnormal{lower}}\lambda_{\textnormal{min}}(\Sigma)^{1/2}}=:\beta_{\textnormal{upper}},$
completing the proof. ∎
Appendix M Monotonicity
Lemma M.1.
For some set $\mathcal{X}$ and $f,g:\mathcal{X}\to\mathbb{R}$, let $x(c)=\arg\min_{x\in\mathcal{X}}(f(x)+cg(x))$, where $f,g$ are such that for some interval $\mathcal{I}\subset\mathbb{R}$, the minimum is attained for all $c\in\mathcal{I}$. Then $G:\mathcal{I}\to\mathbb{R}$, $G:c\mapsto g(x(c))$ is non-increasing in $c$.
Proof.
Let $c_{1},c_{2}\in\mathcal{I}$, $c_{1}<c_{2}$. At $c=c_{1}$, the minimizer $x(c_{1})$ satisfies:
$f(x(c_{1}))+c_{1}g(x(c_{1}))\leq f(x(c_{2}))+c_{1}g(x(c_{2})).$
At $c=c_{2}$, the minimizer $x(c_{2})$ satisfies:
$f(x(c_{2}))+c_{2}g(x(c_{2}))\leq f(x(c_{1}))+c_{2}g(x(c_{1})).$
Adding the two inequalities, we find
$\big{[}f(x(c_{1}))+c_{1}g(x(c_{1}))\big{]}+\big{[}f(x(c_{2}))+c_{2}g(x(c_{2}))\big{]}\leq\big{[}f(x(c_{1}))+c_{2}g(x(c_{1}))\big{]}+\big{[}f(x(c_{2}))+c_{1}g(x(c_{2}))\big{]}.$
Subtracting the common terms $f(x(c_{1}))+f(x(c_{2}))$ leads to
$c_{1}g(x(c_{1}))+c_{2}g(x(c_{2}))\leq c_{2}g(x(c_{1}))+c_{1}g(x(c_{2})).$
Rearranging, and factoring out $c_{1}$ and $c_{2}$, we find
$c_{1}\big{[}g(x(c_{1}))-g(x(c_{2}))\big{]}-c_{2}\big{[}g(x(c_{1}))-g(x(c_{2}))\big{]}\leq 0.$
Thus, $(c_{1}-c_{2})\big{[}g(x(c_{1}))-g(x(c_{2}))\big{]}\leq 0.$ Since $c_{2}-c_{1}>0$, the inequality implies $g(x(c_{1}))\geq g(x(c_{2}))$, as desired. ∎
Appendix N Helper lemmas
Lemma N.1.
If $\alpha\leqslant 0.5$, then $\alpha|c-s|\leqslant\ell_{\alpha}(c,s)\leqslant(1-\alpha)|c-s|$ for all $c,s\in\mathbb{R}$.
Proof.
If $s\geqslant c$, then $\ell_{\alpha}(c,s)=(1-\alpha)(s-c)$. Since $s-c\geqslant 0$ and $\alpha\leqslant 1-\alpha$, we have $\alpha(s-c)\leqslant\ell_{\alpha}(c,s)\leqslant(1-\alpha)(s-c)$, which implies $\alpha|c-s|\leqslant\ell_{\alpha}(c,s)\leqslant(1-\alpha)|c-s|$. If $s<c$, then $\ell_{\alpha}(c,s)=\alpha(c-s)$. Since $c-s>0$ and $\alpha\leqslant 1-\alpha$, we have $\alpha(c-s)\leqslant\ell_{\alpha}(c,s)\leqslant(1-\alpha)(c-s)$, which implies $\alpha|c-s|\leqslant\ell_{\alpha}(c,s)\leqslant(1-\alpha)|c-s|$. ∎
Lemma N.2.
If $\alpha\leqslant 0.5$, then the map $\mathbb{R}\to\mathbb{R}$ given by $c\mapsto\ell_{\alpha}(c,s)$ is $(1-\alpha)$-Lipschitz.
Proof.
If $s\leqslant c_{1}\leqslant c_{2}$, we have $0\leqslant\ell_{\alpha}(c_{2},s)-\ell_{\alpha}(c_{1},s)=\alpha(c_{2}-c_{1})$, which by $\alpha\leqslant 0.5$ is at most $(1-\alpha)(c_{2}-c_{1})$. Hence $|\ell_{\alpha}(c_{2},s)-\ell_{\alpha}(c_{1},s)|\leqslant(1-\alpha)|c_{2}-c_{1}|$. If $c_{1}\leqslant s\leqslant c_{2}$ and $\ell_{\alpha}(c_{2},s)\geqslant\ell_{\alpha}(c_{1},s)$, then we have
$0\leqslant\ell_{\alpha}(c_{2},s)-\ell_{\alpha}(c_{1},s)=\alpha(c_{2}-s)-(1-\alpha)(s-c_{1})\leqslant\alpha(c_{2}-s)+\alpha(s-c_{1})=\alpha(c_{2}-c_{1}),$
which by $\alpha\leqslant 0.5$ implies $|\ell_{\alpha}(c_{2},s)-\ell_{\alpha}(c_{1},s)|\leqslant(1-\alpha)|c_{2}-c_{1}|$. If $c_{1}\leqslant s\leqslant c_{2}$ and $\ell_{\alpha}(c_{2},s)\leqslant\ell_{\alpha}(c_{1},s)$, then
$0\leqslant\ell_{\alpha}(c_{1},s)-\ell_{\alpha}(c_{2},s)=(1-\alpha)(s-c_{1})-\alpha(c_{2}-s)\leqslant(1-\alpha)(s-c_{1})+(1-\alpha)(c_{2}-s)=(1-\alpha)(c_{2}-c_{1}),$
hence $|\ell_{\alpha}(c_{2},s)-\ell_{\alpha}(c_{1},s)|\leqslant(1-\alpha)|c_{2}-c_{1}|$. Finally, if $c_{1}\leqslant c_{2}\leqslant s$, we have $0\leqslant\ell_{\alpha}(c_{1},s)-\ell_{\alpha}(c_{2},s)=(1-\alpha)(c_{2}-c_{1})$, hence $|\ell_{\alpha}(c_{2},s)-\ell_{\alpha}(c_{1},s)|\leqslant(1-\alpha)|c_{2}-c_{1}|$. ∎
Lemma N.3.
The map $\mathcal{H}\to\mathbb{R}$ given by $h\mapsto\ell_{\alpha}(h(x),s)$ is convex for all $x\in\mathcal{X}$ and $s\in\mathbb{R}$.
Proof.
Write $h(x)=\langle\gamma,\Phi\rangle$ for $\gamma\in\mathbb{R}^{d}$. It suffices to show that the mapping $\mathbb{R}^{d}\to\mathbb{R}$ given by $\gamma\mapsto\ell_{\alpha}(\gamma^{\top}\Phi(x),s)$ is convex. But this map is the composition of the linear function $\mathbb{R}^{d}\to\mathbb{R}$ given by $\gamma\mapsto\gamma^{\top}\Phi(x)$ and the convex function $\mathbb{R}\to\mathbb{R}$ given by $c\mapsto\ell_{\alpha}(c,s)$, hence it is convex. ∎
Lemma N.4.
Under 1, if $h\in\mathcal{H}$, then $\sup_{x\in\mathcal{X}}|h(x)|\leqslant C_{\Phi}\|h\|$, where we use the norm given by $\|h\|=\|\gamma\|_{2}$ for $h=\langle\gamma,\Phi\rangle$. In particular, if $h\in\mathcal{H}_{B}$, then $\sup_{x\in\mathcal{X}}|h(x)|\leqslant BC_{\Phi}$.
Proof.
Writing $h=\langle\gamma,\Phi\rangle$ for $\gamma\in\mathbb{R}^{d}$, we have $\sup_{x\in\mathcal{X}}|h(x)|=\sup_{x\in\mathcal{X}}|\langle\gamma,\Phi(x)\rangle|\leqslant\sup_{x\in\mathcal{X}}\|\gamma\|_{2}\|\Phi(x)\|_{2}\leqslant C_{\Phi}\|h\|$, where in the second step we applied the Cauchy-Schwarz inequality. ∎
Lemma N.5.
Consider the function $\varphi:\mathbb{R}^{d}\to\mathbb{R}$ given by $\varphi(\gamma)=\mathbb{E}_{1}[\ell_{\alpha}(h_{\gamma}(X),S)]$, where $h:=h_{\gamma}:\mathcal{X}\to\mathbb{R}$ is given by $h(x)=\langle\gamma,\Phi(x)\rangle$ for all $x\in\mathcal{X}$. Then under 1 and 5, $\varphi$ is twice-differentiable, with gradient and Hessian given by
$\nabla_{\gamma}\varphi(\gamma)=\mathbb{E}_{1}[(\mathbb{P}_{S|X}[h(X)>S]-(1-\alpha))\Phi(X)],\qquad\nabla_{\gamma}^{2}\varphi(\gamma)=\mathbb{E}_{1}[f_{S|X}(h(X))\Phi(X)\Phi(X)^{\top}].$
Consequently, given $\tilde{\gamma}\in\mathbb{R}^{d}$, defining $g:\mathcal{X}\to\mathbb{R}$ as $g(x)=\langle\tilde{\gamma},\Phi(x)\rangle$ for all $x\in\mathcal{X}$, the directional derivative of $\varphi:\mathcal{H}\to\mathbb{R}$ in the direction $g$ is given by $\langle\tilde{\gamma},\nabla_{\gamma}\varphi(\gamma)\rangle=\mathbb{E}_{1}[(\mathbb{P}_{S|X}[h(X)>S]-(1-\alpha))g(X)]$.
Proof.
For each $x\in\mathcal{X}$, define the function $\eta(\cdot;x):\mathbb{R}\to\mathbb{R}$ given, for all $u$, by $\eta(u;s)=\mathbb{E}_{S|X=x}[\ell_{\alpha}(u,S)]$. For each $s\in\mathbb{R}$, define the function $\chi(\cdot;s):\mathbb{R}\to\mathbb{R}$, where for all $u$, $\chi(u;s)=\alpha\mathbf{1}[u>s]-(1-\alpha)\mathbf{1}[u\leqslant s]$.
By the definition of the pinball loss $\ell_{\alpha}(\cdot,\cdot)$, and since by 5 the conditional density $f_{S|X=x}(\cdot)$ of $S|X=x$ exists for all $x\in\mathcal{X}$, the derivative of $\ell_{\alpha}(u,S)$ with respect to $u$ agrees with the random variable $\chi(u;S)$ almost surely with respect to the distribution $S|X=x$. Also, note that for fixed $u\in\mathbb{R}$, $|\chi(u;S)|$ is bounded by the constant $(1-\alpha)$. By the dominated convergence theorem, it follows that $u\mapsto\eta(u;x)$ is differentiable, and that its derivative equals $\frac{\partial}{\partial u}\eta(u;x)=\mathbb{E}_{S|X=x}[\chi(u;S)]$, which, by the formula for $\chi(u;S)$, can be written as $\alpha\mathbb{P}_{S|X=x}[u>S]-(1-\alpha)\mathbb{P}_{S|X=x}[u\leqslant S]$. Thus for all $u\in\mathbb{R}$ and $x\in\mathcal{X},$ we may write $\frac{\partial}{\partial u}\eta(u;x)=\mathbb{P}_{S|X=x}[u>S]-(1-\alpha)$. Since by 5 the conditional density $f_{S|X=x}$ of the distribution $S|X=x$ exists for all $x\in\mathcal{X}$, it follows that the cdf $u\mapsto\mathbb{P}_{S|X=x}[u>S]$ is differentiable for all $u\in\mathbb{R}$ and all $x\in\mathcal{X}$ with derivative given by $u\mapsto f_{S|X=x}(u)$. Thus the map $u\mapsto\frac{\partial}{\partial u}\eta(u;x)$ is differentiable for all $x\in\mathcal{X}$ with derivative given by $u\mapsto f_{S|X=x}(u)$. In particular, $\eta(\cdot;x)$ is twice-differentiable with second derivative given by $f_{S|X=x}(\cdot)$.
Next, for each $x\in\mathcal{X}$, define the function $\psi(\cdot;x):\mathbb{R}^{d}\to\mathbb{R}$ given by $\psi(\gamma;x)=\mathbb{E}_{S|X=x}[\ell_{\alpha}(h_{\gamma}(x),S)]$, where $h=h_{\gamma}=\langle\gamma,\Phi\rangle$. For each $x\in\mathcal{X}$, let $\textnormal{ev}(\cdot;x):\mathbb{R}^{d}\mapsto\mathbb{R}$ be given by $\textnormal{ev}(\gamma;x)=h_{\gamma}(x)$, where $h=h_{\gamma}=\langle\gamma,\Phi\rangle$. Then $\psi(\cdot;x)$ is given by the composition $\eta(\cdot;x)\circ\textnormal{ev}(\cdot;x)$. Since $\textnormal{ev}(\gamma;x)=\langle\gamma,\Phi(x)\rangle$, $\textnormal{ev}(\cdot;x)$ is linear, it is smooth. Its gradient is given by $\nabla_{\gamma}\textnormal{ev}(\gamma;x)=\Phi(x)$ for all $\gamma\in\mathbb{R}^{d}$, and its Hessian is zero. It follows that $\psi(\cdot;x)$ is twice-differentiable. By the chain rule, the gradient of $\psi(\cdot;x)$ is given by
$\nabla_{\gamma}\psi(\gamma;x)=\frac{\partial}{\partial u}\eta(u;x)\bigg{|}_{u=\textnormal{ev}(\gamma;x)}\cdot\nabla_{\gamma}\textnormal{ev}(\gamma;x)=(\mathbb{P}_{S|X=x}[h(x)>S]-(1-\alpha))\Phi(x).$
Since the map $\gamma\mapsto\mathbb{P}_{S|X=x}[h(x)>S]-(1-\alpha)$ is given by the composition $\frac{\partial}{\partial u}\eta(\cdot;x)\circ\textnormal{ev}(\cdot;x)$, we may again apply the chain rule to deduce that the Hessian of $\psi(\cdot;x)$ is given by
$\nabla_{\gamma}^{2}\psi(\gamma;x)=\frac{\partial^{2}}{\partial u^{2}}\eta(u;x)\bigg{|}_{u=\textnormal{ev}(\gamma;x)}\cdot\nabla_{\gamma}\textnormal{ev}(\gamma;x)\cdot\Phi(x)^{\top}=f_{S|X=x}(h(x))\Phi(x)\Phi(x)^{\top}.$
Returning to our original function $\varphi$, note that by the tower property, $\varphi(\gamma)=\mathbb{E}_{1}[\psi(\gamma;X)]$. Note that $\|\nabla_{\gamma}\psi(\gamma;x)\|_{2}$ is at most
$|\mathbb{P}_{S|X=x}[h(x)>S]-(1-\alpha)|\|\Phi(x)\|_{2}\leqslant(|\mathbb{P}_{S|X=x}[h(x)>S]|+(1-\alpha))\|\Phi(x)\|_{2}\leqslant(2-\alpha)C_{\Phi},$
where in the first step we used the triangle inequality, and in the second step we used the fact that $\mathbb{P}_{S|X=x}[h(x)>S]\leqslant 1$ and 1. Similarly, we may bound the Frobenius norm $\|\cdot\|_{F}$ of $\nabla_{\gamma}^{2}\psi(\gamma;x)$ by
$|f_{S|X=x}(h(x))|\|\Phi(x)\Phi(x)^{\top}\|_{F}\leqslant C_{f}\|\Phi(x)\|_{2}^{2}\leqslant C_{f}C_{\Phi}^{2},$
where in the first step we used 1, the identity $\|vv^{\top}\|_{F}=\|v\|_{2}^{2}$, and in the second step we used 5. Since the entries of $\nabla_{\gamma}\psi(\cdot;x)$ and $\nabla_{\gamma}^{2}\psi(\cdot;x)$ are bounded by constants, we may apply the dominated convergence theorem to deduce that $\varphi$ is twice-differentiable, with gradient given by $\nabla_{\gamma}\varphi(\gamma)=\mathbb{E}_{1}[\nabla_{\gamma}\psi(\gamma;X)]$ and Hessian given by $\nabla_{\gamma}^{2}\varphi(\gamma)=\mathbb{E}_{1}[\nabla_{\gamma}^{2}\psi(\gamma;X)]$.
Finally, since the directional derivative of $\varphi$ in the direction $g$ is defined as $\langle\tilde{\gamma},\nabla_{\gamma}\varphi(\gamma)\rangle$, we may plug in our expression for the gradient to deduce
$\langle\tilde{\gamma},\nabla_{\gamma}\varphi(\gamma)\rangle$ $=\langle\tilde{\gamma},\mathbb{E}_{1}[(\mathbb{P}_{S|X}[h(X)>S]-(1-\alpha))\Phi(X)]\rangle$
$=\mathbb{E}_{1}[(\mathbb{P}_{S|X}[h(X)>S]-(1-\alpha))\langle\tilde{\gamma},\Phi(X)\rangle]=\mathbb{E}_{1}[(\mathbb{P}_{S|X}[h(X)>S]-(1-\alpha))g(X)].$
The result follows. ∎